### v1

# Improving governance outcomes through AI documentation: Bridging theory and practice

Abstract.
Documentation plays a crucial role in both external accountability and internal governance of AI systems. Although there are many proposals for documenting AI data, models, systems, and methods, the ways these practices enhance governance as well as the challenges practitioners and organizations face with documentation remain underexplored. In this paper, we analyze 37 proposed documentation frameworks and 21 empirical studies evaluating their use. We identify potential hypotheses about how documentation can strengthen governance—such as informing stakeholders about AI risks and usage, fostering collaboration, encouraging ethical reflection, and reinforcing best practices. However, empirical evidence shows that practitioners often encounter obstacles that prevent documentation from achieving these goals. We also highlight key considerations for organizations when designing documentation, such as determining the appropriate level of detail and balancing automation in the process. Finally, we offer recommendations for further research and for implementing effective documentation practices in real-world contexts.
1. Introduction
Transparency in AI is widely regarded as essential for calibrating trust and supporting accountability (Liao and Vaughan, 2023). By providing stakeholders with a clear view of an AI system’s composition, operation, and development process, transparency allows for informed oversight and critical evaluation. When AI systems are thoroughly documented, these documentation artifacts offer invaluable insights into important system features such as training data, algorithms, and risk management strategies. This information can support two broad groups of external stakeholders. Documentation can help downstream deployers and users understand how the system functions and what risks it conveys. It can also assist policymakers and researchers in holding companies accountable for the negative consequences of their AI technologies.
Yet the importance of documentation extends beyond external accountability.111Though the concepts of documentation and transparency are often blurred, complete documentation about an AI system does not guarantee that information will be made available to interested stakeholders, nor do transparency artifacts always reflect the full or most relevant details of an AI system even though such details may have been captured in internal documents. We distinguish between the concepts in this report in order to provide greater clarity about the presumptive goals of documentation and an exploration of what practices may more likely support those objectives Internally, documentation serves as a critical tool for managing AI systems throughout their lifecycle for a wide range of stakeholders. For example, documentation could help downstream technical practitioners understand the strengths, limitations, and risks of training a model on a given dataset. It might also help compliance and governance teams assess a potential system use case’s compliance with company policies or legal requirements. Or the process of producing documentation might lead practitioners to make different decisions about how to sufficiently mitigate potential harms. As such, documentation is a foundational building block of effective AI governance.
Despite its clear benefits, creating effective documentation is a complex task. No single form of documentation can meet the diverse needs of all stakeholder groups equally. For example, while detailed technical reports like those for Llama-2(Touvron et al., 2023) and GPT-4(OpenAI, 2023), which can exceed 50 pages and contain dense technical language, are valuable for AI researchers, they may not be accessible to non-technical stakeholders. Similarly, documentation intended to satisfy compliance requirements might fall short in providing practical guidance for those deploying AI systems.
To maximize the impact of documentation on AI governance, organizations must carefully define their goals and identify who will produce, maintain, and use the documentation. Tailoring the scope, level of detail, and format of documentation to suit the intended audiences and purposes is crucial. Effective documentation design requires balancing the needs of both the documentation process and the resulting artifacts, ensuring that they collectively support the organization’s governance objectives.
In this paper, we explore how documentation processes and artifacts can best support AI governance and risk management goals. Our research draws from a qualitative analysis of academic and gray literature,222Gray literature refers to publications that have not been peer-reviewed, but that present detailed theories or research findings including 37 proposed documentation frameworks and 21 papers with relevant empirical findings. From this analysis, we identify four mechanisms through which documentation is hypothesized to improve AI governance:
Informing stakeholders about responsible use,
Facilitating collaboration on AI risks,
Encouraging ethical deliberation, and
Improving overall governance and development practices.
We also examine the practical barriers that may hinder the effectiveness of these mechanisms and the considerations or trade-offs organizations face when designing and implementing documentation practices. These considerations include balancing customization with standardization, deciding between single or multiple forms of documentation, determining the extent of detail to include, choosing the appropriate level of automation to employ in the documentation process, and selecting between static or interactive formats. Based on our analysis, we offer recommendations to help organizations develop documentation strategies that align with their governance goals and stakeholder needs.
The primary contribution of our research is that our synthesis identifies patterns, gaps, and emerging trends across individual publications about documentation approaches and evaluations. By offering a holistic analysis of how AI documentation can support robust governance in real-world contexts, we clarify which documentation practices are likely to be most effective, under what conditions, and for which stakeholders. As such, our research can serves as a valuable resource for both researchers and practitioners seeking to refine and enhance their approaches to AI documentation in applied AI development contexts.
2. What organizations could document about AI systems
Documenting the various components of AI systems—data, models, systems, and methods—is essential for effective governance and informed decision-making. This documentation not only provides insight into the development process but also helps internal and external stakeholders understand the characteristics and potential risks associated with AI systems.
Data is a fundamental component of AI systems since it significantly influences model behavior and performance. When models are trained on datasets that don’t align with their deployment contexts or that contain biases, the outcomes can be problematic, reinforcing those biases or leading to poor performance (Gebru et al., 2021). Data documentation, which might include details about dataset characteristics, collection methods, preprocessing, and known limitations, allows practitioners to identify potential issues early on, supporting better model performance and informing the decisions of those who use the documentation (Bender and Friedman, 2018; Chmielinski et al., 2020; Díaz et al., 2023; Gebru et al., 2021; Marone and Van Durme, 2023; Miceli et al., 2021; Papakyriakopoulos et al., 2023; Pushkarna et al., 2022; Roman et al., 2023; Rostamzadeh et al., 2022; Soh, 2021; Stoyanovich and Howe, 2019; Subramaniam et al., 2023; Sun et al., 2019; Zheng et al., 2022). By tailoring documentation to the specific data type and application context, organizations can enhance its utility, drawing on established data documentation frameworks to guide this process.
Model documentation also enhances AI development and governance, providing a window into a model’s creation, performance, and risks (Crisan et al., 2022; McMillan-Major et al., 2021; Mitchell et al., 2019; Shen et al., 2022; Stoyanovich and Howe, 2019). Documenting aspects such as the model’s architecture, training procedures, intended use cases, and evaluation methods enables practitioners to assess potential risks and responsibly adapt models to specific contexts. For models developed in stages—such as those pretrained on large datasets and later fine-tuned—documenting each stage is essential. This helps downstream users understand the development pipeline and its implications for their intended use.
AI services or applications in the real world typically consist of complex systems rather than isolated models. These systems often include multiple components, such as machine learning models, trust and safety filters, third-party software integrations, and user interfaces. The interaction among these components significantly affects the system’s overall performance and risk profile — especially because the properties or behavior of individual components may change when integrated into a system (Smart et al., 2024). Recognizing that AI systems are more than just the sum of their parts, researchers have suggested that organizations should document not only each individual component but also the system as a whole (Arnold et al., 2019; Baracaldo et al., 2022; Brajovic et al., 2023; Green et al., 2021; Procope et al., 2022; Yang et al., 2018; Blasch et al., 2021). System documentation could provide an overview, including objectives, inputs, outputs, and a diagram illustrating how the components interact. This comprehensive approach ensures that practitioners have a clear understanding of the system’s behavior in its entirety.
Organizations may also find it important to document the processes and procedures that shape the development and deployment of AI systems and their components. This documentation could include how humans annotated or evaluated training or evaluation data (Shimorina and Belz, 2021), any privacy, legal, and ethical reviews or audits conducted (Raji and Yang, 2020), and any risk mitigations that were applied. Investing time in this type of documentation may be especially beneficial because it cannot be deduced merely from examining the system components. If such meta-data about the process is not documented throughout the development lifecycle, it may be very difficult to reconstruct later (Belz et al., 2023; Reid and Williams, 2023). As a result, it is ideal to produce documentation on development processes concurrently with the development of the system.
Organizations may also benefit from documenting machine learning methods (Adkins et al., 2022; Baracaldo et al., 2022; Sokol and Flach, 2020; Tagliabue et al., 2021), such as techniques applied in computer vision applications (Adkins et al., 2022) and strategies for enhancing model explainability (Sokol and Flach, 2020). Documentation of the application of the AI model or system to specific AI tasks or use cases (Hupont and Gomez, 2022; Mohammad, 2022) can also help practitioners understand the benefits and limitations of these applications. Information about methods and processes can provide valuable insights into what potential approaches a team might take to develop an AI system. Such insights can help practitioners determine when particular approaches are more or less appropriate and what risks different approaches may confer.
Within companies, many stakeholders may contribute to and use documentation (Micheli et al., 2023). Those working on data collection, annotation, and curation might contribute to data documentation, while engineers and data scientists might document and use information related to data, models, and application infrastructure. UX designers, researchers, and product managers might rely on this documentation to inform product designs and create clear public explanations. Professionals focused on responsible AI, governance, and compliance could use documentation to identify and manage risks. Organizational leaders, in turn, may use these insights to make critical decisions about resource allocation, risk management, and the launch of AI-powered products.
We can categorize documentation stakeholders into two broad groups: documentation producers and documentation consumers. Documentation producers are those who actively contribute to the generation of documentation artifacts, while documentation consumers are those who read and use it. While this grouping simplifies some important distinctions, it serves as a helpful heuristic for understanding the opportunities and tensions between groups with both shared and distinct goals. We consider both producers and consumers and their interplay in our analysis.
We also note that ”AI documentation” could refer to either the process of producing documentation or the resulting artifacts. While documentation artifacts inform stakeholders about the responsible use of AI, the process of creating these documents can also institutionalize best practices and foster a culture of risk management.. Therefore, when considering how documentation can improve governance outcomes, we address both artifacts and processes.
3. Methods
3.1. Publication Sampling
We employed a purposeful sampling approach (Palinkas et al., 2015) to identify proposed documentation frameworks and relevant empirical studies. Given our specific interest in how AI documentation can support governance, our use of a purposeful sampling approach allowed us to focus on proposed frameworks and empirical studies that are directly relevant to our research questions. It also afforded us the flexibility to iteratively refine our sample of papers based on emerging findings.
We identified an initial seed sample of publications from a systematic review focused on AI documentation methods relevant to EU regulation (Micheli et al., 2023). We cross-referenced this initial sample against the references used by the Partnership on AI in developing their ABOUT ML framework,333https://partnershiponai.org/workstream/about-ml/ which seeks to provide technology organizations with guidance on what to document about their AI systems. We chose these initial sources because they allowed us to focus on publications that aimed to establish best practices for industry documentation and address policy initiatives geared toward institutionalizing these practices. We reviewed the cited references within each of these initial works for additional proposed documentation frameworks and empirical studies related to documentation. We also consulted with academic researchers and industry practitioners with experience producing and using documentation to identify additional frameworks and studies. We excluded papers that centered on tools for implementing documentation (e.g., automated documentation code libraries) or that were not specifically focused on AI system documentation, such as fairness checklists.
This process yielded a sample of 37 proposed frameworks, eight of which included empirical evidence related to their implementation. Because of the relative dearth of empirical findings within publications proposing approaches, we further expanded our sample by searching for references within these 37 approaches that empirically evaluated documentation methods. We also conducted searches in the archives of the ACM Conference on Human Factors in Computing Systems and the ACM Conference on Computer Supported Cooperative Work using terms such as ”datasheet,” ”model card,” ”AI documentation,” ”model documentation,” ”data card,” and ”data documentation.” Together, these methods allowed us to identify 13 additional publications with empirical findings.
Recognizing the importance of applied perspectives, we included non-peer-reviewed papers, such as technical whitepapers published by industry actors, acknowledging that they often provide insights not found in academic literature. While our sampling approach was not exhaustive, we reached theoretical saturation after analyzing the selected publications, as no new themes emerged. Therefore, we concluded our sampling at this point.
3.2. Data Analysis
Our data analysis followed an abductive approach (Timmermans and Tavory, 2012; Tavory and Timmermans, 2014), iterating between inductively derived codes, as in grounded theory methods (Strauss and Corbin, 1990; Charmaz, 2014), and codes based on the theoretical motivations of our research. The first author conducted the initial coding by applying descriptive codes aligned with our research goals, such as identifying how AI documentation enhances internal governance and the challenges of using documentation as a governance tool. After discussing the results with the second author, we refined these codes into higher-order categories during the axial coding phase. For instance, codes like ”identify intended purpose” and ”specify out-of-scope uses” were grouped under the category ”inform downstream documentation consumers.”
In the final thematic coding phase, we grouped axial codes into broader theoretical mechanisms through which AI documentation could support robust governance, the challenges in achieving these impacts, and the design trade-offs faced by organizations.
4. Results
4.1. Publication Sample
The publications included in our analysis are listed in Table LABEL:tab:ai_docs. For each framework, we classified whether the approach primarily focused on data; models; systems; or methods, tasks, and processes. When a framework did not fit neatly into one of these categories, we assigned it to the most appropriate category or categories based on its primary focus.
We also categorized the type of evaluation each framework or independent empirical research study employed. Frameworks employing a ”feasibility analysis” are those where the framework’s authors or another group applied the framework to create documentation for a hypothetical or actual dataset, model, system, or method. This type of analysis demonstrates that the framework could theoretically be used for its intended purpose but does not involve empirical evaluation with practitioners in research or real-world settings. If a study developed a documentation artifact for the purpose of an empirical study, we classified this as part of the empirical study rather than as a feasibility analysis, as empirical studies offer a more rigorous evaluation.
We classify publications as employing a ”practitioner lab study” when the evaluation involved practitioners within a controlled research setting. We classify ”practitioner real-world studies” as those examining practitioner methods and practices within their real-world work environments. Both lab and real-world studies have unique strengths, and neither is inherently more rigorous or useful than the other. We define ”artifact studies” as studies of publicly available documentation artifacts such as Github repository documentation or Hugging Face model cards.
In some instances, framework authors mentioned consulting relevant stakeholders during the design or refinement of their framework. However, if these consultations are only briefly mentioned or not elaborated upon, we do not classify the work as including a practitioner study.
| Num. | Author | Year | Framework Type | Evaluation Type |
| 1 | Adkins et al. | 2022 | Method, process, or task | Feasibility analysis |
| 2 | Arnold et al. | 2019 | System | Feasibility analysis |
| 3 | Baracaldo et al. | 2022 | System | Feasibility analysis |
| 4 | Bender & Friedman | 2018 | Data | Feasibility analysis |
| 5 | Bhat et al. | 2023 | Evaluation only | Practitioner lab study |
| 6 | Blasch et al. | 2020 | System | Feasibility analysis |
| 7 | Boyd | 2021 | Evaluation only | Practitioner lab study |
| 8 | Brajovic et al. | 2023 | System | Feasibility analysis |
| 9 | Chang & Custis | 2022 | Evaluation only | Practitioner real-world study |
| 10 | Chmielinski et al. | 2020 | Data | Feasibility analysis |
| 11 | Chmielinski et al. | 2024 | Method, process, or task | None |
| 12 | Crisan et al. | 2022 | Model | Feasibility analysis, Practitioner lab study |
| 13 | Díaz et al. | 2023 | Data | Feasibility analysis |
| 14 | Gebru et al. | 2021 | Data | Feasibility analysis |
| 15 | Geiger et al. | 2020 | Evaluation only | Artifact study |
| 16 | Gilbert et al. | 2023 | System | Feasibility analysis |
| 17 | Heger et al. | 2022 | Evaluation only | Practitioner real-world study |
| 18 | Hind et al. | 2019 | Evaluation only | Practitioner real-world study |
| 19 | Holland et al. | 2018 | Data | Feasibility analysis |
| 20 | Hupont & Gomez | 2022 | Method, process, or task | Feasibility analysis |
| 21 | Liang et al. | 2024 | Evaluation only | Artifact study |
| 22 | Liao et al. | 2023 | Evaluation only | Practitioner lab study |
| 23 | Marone & Van Durme | 2023 | Data | Feasibility analysis |
| 24 | McMillan-Major et al. | 2021 | Data, Model | Feasibility analysis |
| 25 | McMillan-Major, Bender, & Friedman | 2024 | Data | Practitioner lab study |
| 26 | Miceli et al. | 2021 | Evaluation only | Practitioner real-world study |
| 27 | Mitchell et al. | 2019 | Model | Feasibility analysis |
| 28 | Mohammad | 2022 | Method, process, or task | Feasibility analysis |
| 29 | Moore, LIao, & Subramonyam | 2023 | Evaluation only | Practitioner lab study |
| 30 | Nunes et al. | 2022 | Evaluation only | Practitioner lab study |
| 31 | Papakyriakopoulos et al. | 2023 | Data | Feasibility analysis |
| 32 | Pepe et al. | 2024 | Evaluation only | Artifact study |
| 33 | Procope et al. | 2022 | System | Feasibility analysis |
| 34 | Pushkarna et al. | 2022 | Data | Practitioner real-world study |
| 35 | Raji & Yang | 2020 | System | None |
| 36 | Reid & Williams | 2023 | Evaluation only | Practitioner real-world study, Artifact study |
| 37 | Richards et al. | 2020 | System | Feasibility analysis |
| 38 | Roman et al. | 2023 | Data | Feasibility analysis |
| 39 | Rostamzadeh et al. | 2022 | Data | Practitioner lab study |
| 40 | Shen et al. | 2022 | Model | Practitioner lab study |
| 41 | Shimorina & Belz | 2021 | Method, process, or task | None |
| 42 | Soh | 2021 | Data | None |
| 43 | Sokol & Flach | 2020 | Method, process, or task | Feasibility analysis |
| 44 | Srinivasan et al. | 2021 | Data | Feasibility analysis |
| 45 | Stoyanovich & Howe | 2019 | Data, Model | Feasibility analysis |
| 46 | Subramaniam et al. | 2023 | Data | Practitioner lab study |
| 47 | Sun et al. | 2019 | Data | Feasibility analysis |
| 48 | Tagliabue et al. | 2021 | Method, process, or task | Feasibility analysis |
| 49 | Yang et al. | 2018 | Method, process, or task | Feasibility analysis |
| 50 | Zheng et al., | 2022 | Data | Practitioner lab study |
4.2. Potential Impacts & Challenges
While empirical evidence directly linking documentation artifacts or processes to improved governance outcomes is limited, the theories proposed in existing documentation frameworks offer compelling hypotheses about how documentation can positively impact governance. Empirical studies involving practitioners, alongside analyses of publicly available documentation artifacts, provide an initial glimpse into the factors that contribute to accurate, high-quality, and comprehensive documentation, as well as the reasons why documentation efforts may sometimes fall short of their goals.
Our analysis identified four possible ways in which documentation could improve governance: informing downstream users about system development and associated risks, encouraging ethical reflection among practitioners, facilitating communication among stakeholders, and enhancing AI development and governance overall. In the following sections, we explore each of these hypothesized impacts in detail and examine the challenges real-world organizations face in realizing these benefits.
4.2.1. Potential impact #1: Documentation could inform downstream stakeholders about effective and responsible use.
The most straightforward potential impact of documentation is that it can provide essential context, such as the original motivation for developing a dataset, model, or system, their intended use cases, and information on system properties necessary for effective and responsible deployment.
Hypotheses. Documentation could inform downstream stakeholders by clearly articulating the intended uses of data, models, and systems and provide guidance on responsible deployment. Practitioners typically develop systems with specific use cases in mind, and other development teams need to understand these intended use cases to assess alignment with their goals. If aligned, teams can then modify the systems to meet their current objectives. Practitioners argue documentation can also clarify underlying assumptions, such as data collection methods and representativeness, which are critical for assessing the suitability of a dataset for specific purposes (Reid and Williams, 2023; Rostamzadeh et al., 2022; Zheng et al., 2022). Also, it can help identify potential deployment issues, such as mismatches between the original training data and the intended deployment context (Boyd, 2021).
Moreover, documentation could assist downstream practitioners in making implementation decisions that support effective and responsible product development. For example, documentation could help practitioners find and compare system components aligned with their goals, including alternatives to AI approaches (Heger et al., 2022; Reid and Williams, 2023; Shen et al., 2022; Zheng et al., 2022). It might also provide statistical summaries of data (Holland et al., 2018) or or model performance evaluations such as area under the curve (AUC) (Mitchell et al., 2019) to help inform developers’ choices of how to train models or implement appropriate guardrails.
Documentation could also help stakeholders ensure that planned use cases comply with legal constraints, regulatory requirements, and organizational policies. For example, certain datasets may be restricted from use if the data were collected without affirmative consent or contain personally identifiable or copyrighted information. Without this awareness, practitioners might inadvertently train a model on such data, only to later discover that they cannot use the model due to policy violations (Roman et al., 2023). Practitioners see avoiding disciplinary action for non-compliance with company policies as a motivating benefit of documentation (Heger et al., 2022).
Challenges. While practitioners acknowledge that the benefits of documentation could outweigh the costs of creating and maintaining it (Heger et al., 2022), when organizations fail to provide clear incentives for high-quality documentation, practitioners are less likely to prioritize it. This lack of motivation can hinder the realization of documentation’s potential benefits, but particularly its ability to inform downstream users.
One major challenge is the lack of organizational incentives for documentation. Organizations typically do not prioritize documentation unless regulations require it for compliance or clients specifically request it (Chang and Custis, 2022; Miceli et al., 2021; Heger et al., 2022). Practitioners often view producing high-quality documentation as less relevant to their evaluations and promotions than development tasks that directly contribute to product outputs (Chang and Custis, 2022; Heger et al., 2022). This disconnect between the benefits of documentation and practitioner objectives is further exacerbated when the practitioners who benefit most from documentation (i.e., consumers) are not the same as those responsible for producing it (i.e., producers). In organizations where governance processes mandate documentation, a lack of familiarity with AI documentation concepts among developers could potentially magnify tensions between product development and governance teams (Ali et al., 2023; Ahlawat et al., 2024). Moreover, practitioners are often more motivated by compliance than by normative considerations, which could lead to blind spots when novel ethical issues arise outside existing policies (Heger et al., 2022).
Facing time constraints and competing demands, practitioners may deprioritize documentation, sometimes cutting corners to save time (Chang and Custis, 2022; Heger et al., 2022). This can diminish the value of documentation artifacts, as evidenced by studies showing that practitioners sometimes reuse content from existing documentation, even when it pertains to different systems, leading to inadequate or incorrect documentation of their current system (Pushkarna et al., 2022; Heger et al., 2022). Practitioners also sometimes leave questions unanswered rather than seek out the necessary information (Heger et al., 2022) or prioritize recording information that is most valuable to them, potentially omitting critical details for others less familiar with the system (Hind et al., 2019). Furthermore, because practitioners often forget relevant information if they do not record it during development, their reconstruction of key system information after the fact can be time-consuming and error prone (Reid and Williams, 2023).
Studies of publicly available documentation artifacts frequently highlight issues with incomplete or incorrect information, which can mislead downstream stakeholders regarding the characteristics of AI systems or components (Pepe et al., 2024; Liang et al., 2024; Geiger et al., 2020). For example, many publicly available model cards lack details about out-of-scope uses, limitations, or environmental impacts (Pepe et al., 2024; Liang et al., 2024). Less than 30% of model cards contain evaluation results, a critical piece of information for downstream users (Liang et al., 2024). Also, some publicly available documentation contains incorrect license information, potentially leading users to inadvertently violate license terms (Pepe et al., 2024).
4.2.2. Potential impact #2: Documentation could support cross-functional collaboration and communication about AI risks.
Documentation has the potential to act as a bridge between different stakeholder groups involved in the AI development lifecycle, including machine learning engineers, data scientists, product managers, user experience designers, researchers, and legal teams. Since no single group oversees all stages of development, effective documentation can facilitate both indirect and direct communication among these diverse teams. Documentation artifacts can convey essential information across organizational boundaries, helping stakeholders who may not interact regularly to align their understanding of the system. Documentation can also prompt and support conversations among stakeholders directly, enabling them to work towards common goals despite their diverse backgrounds, frames of reference, or areas of expertise. This bridging function of documentation could break down organizational silos(Chmielinski et al., 2024) and promote effective collaboration across different teams (Crisan et al., 2022; Gilbert et al., 2023; Heger et al., 2022; Mohammad, 2022; Pushkarna et al., 2022; Raji and Yang, 2020; Richards et al., 2020; Shen et al., 2022; Srinivasan et al., 2021).
Hypotheses. AI documentation could serve as an initial point of reference for downstream practitioners, offering a basic overview of the system that allows them to build foundational knowledge before engaging with developers for more in-depth discussions. Rather than acting as a comprehensive, self-contained repository of information, documentation artifacts could provide an initial layer of understanding, enabling downstream practitioners to identify relevant areas for further exploration and formulate more informed questions when collaborating with developers. This could foster more productive and targeted interactions between teams, enhancing their ability to work with or govern AI systems efficiently. Empirical research suggests that practitioners often approach documentation this way. For instance, UX professionals often prefer direct discussions with data scientists rather than relying solely on documentation artifacts (Liao et al., 2023). Similarly, many AI practitioners review existing documentation only briefly before seeking further clarification through meetings or discussions (Heger et al., 2022).
In more robust forms, documentation can facilitate deeper deliberation among stakeholders about the system’s design and the documentation process itself. For example, documentation has been shown to aid teams in collectively reflecting on socially constructed concepts, such as gender, and making informed decisions about how to annotate such attributes in datasets (Pushkarna et al., 2022). Also, the process of creating documentation can help stakeholders identify and discuss trade-offs between competing objectives in system design (Shen et al., 2022).
By having some stakeholders define requirements for documentation artifacts, others generate content, and still others assess the quality of the artifact, the documentation process can help stakeholders understand each others’ needs and constraints in ways that could promote quality, efficiency, and responsibility in the development process (Chang and Custis, 2022). Over time, engaging with documentation could also improve the technical literacy of stakeholders, enabling them to create products that better meet user needs while minimizing harm (Liao et al., 2023).
Challenges. Several challenges complicate the role of documentation in facilitating cross-functional collaboration. The modern AI supply chain, characterized by the development of general-purpose models that can be adapted to a variety of applications downstream (Vaswani et al., 2017), often limits opportunities for direct communication between stakeholders. In some cases, different teams within the same organization may adapt a general-purpose model for various specific tasks. In these instances, documentation might prompt the downstream team to directly reach out to the original development team to discuss the system’s capabilities and limitations. In other cases, a team deploying a model may rely on a model developed by a different organization (Jones, 2023). In this case, it will typically be more difficult if not impossible for the downstream deployment team to directly communicate with the original development team.
In either situation, the original developers may not fully anticipate the variety of downstream uses, leading to documentation that lacks critical details needed to identify potential risks or harms. Particularly when upstream and downstream teams cannot communicate directly, it will be more challenging for them to work together to understand the model’s risks within a given deployment context. This challenge is especially pronounced for less technical downstream stakeholders, who may struggle to understand the functionality and risks associated with these models.
More interactive forms of documentation, such as those that allow users to engage with the models directly or customize the level of detail presented, could help mitigate these issues by making the information more accessible and relevant (Crisan et al., 2022; Liao et al., 2023; Moore et al., 2023), especially when direct communication between teams is limited. However, further research is needed to determine the most effective ways to document general-purpose systems while avoiding overwhelming documentation consumers with excessive or irrelevant details,
Another major challenge is that documentation that is misaligned with the needs of its intended consumers can complicate rather than facilitate stakeholder communication. If documentation artifacts are overly technical or jargony, documentation consumers who are less technical or less familiar with the system may struggle to understand what details are most relevant, or misinterpret described practices (Crisan et al., 2022; Liao et al., 2023; Miceli et al., 2021; Shen et al., 2022); if documentation lacks relevant details, documentation consumers may overlook risks that become apparent only later in the development process, which may require teams to substantially backtrack. In either case, mismatch in comprehension can contribute to needless disruption to normal workflows. If practitioners perceive that documentation leads to seemingly egregious friction due to misunderstandings, they may be more prone to oversimplify the documentation they produce or gloss over relevant details that they worry will lead to confusion.
4.2.3. Potential impact #3: Documentation could prompt practitioners to deliberate and act on the ethical impacts of AI
Researchers have suggested that documentation has the potential to encourage practitioners to consider the ethical implications of their systems, leading to more responsible development and use of data, models, and systems. By integrating ethical reflection into the documentation process, practitioners may become more aware of the potential harms their systems could cause, prompting them to make more informed decisions.
Hypotheses. One hypothesis is that when downstream documentation consumers consult documentation on data, models, and systems, it can prompt consideration of potential harms. This consideration could lead practitioners to make more careful decisions about if and when to use certain systems or components. Yet, empirical support for this hypothesis is inconsistent. One study of 23 practitioners examined whether reviewing data documentation artifacts could encourage practitioners to notice ethical issues, contextualize them, and formulate risk mitigation plans (Boyd, 2021). One subset received a scenario description, dataset, and data documentation, while the rest received only the scenario description and dataset. The researcher found that those with data documentation were more likely to notice ethical issues unprompted, though those without it often recognized concerns when the researcher pressed them — but practitioners had difficulty formulating action plans regardless of whether they had access to documentation. These findings suggest that while documentation artifacts can aid ethical deliberation, documentation is not always necessary and likely not sufficient to enable mitigation of ethical issues.
Documentation also might sensitize producers to the ethical impacts of their work. Documentation frameworks often ask documentation producers to consider and describe the ethical dimensions of their systems. For example, the one framework prompts practitioners to think about privacy implications by inquiring about sensitive information in data (Gebru et al., 2021). Another asks about known biases, ethical issues, and safety concerns (Arnold et al., 2019), and several frameworks encourage reflection on how systems might negatively impact marginalized users or populations (Brajovic et al., 2023; McMillan-Major et al., 2021; Shen et al., 2022).
Some researchers have suggested that documentation might be effective at promoting ethical action, even if it only engages practitioners in critical reflection relatively superficially (Bender and Friedman, 2018). Yet others have taken a more explicit, deliberative approach. Drawing from value-sensitive design (Friedman et al., 2006)–a method that helps technologists identify and understand normative judgments in the development process–Shen and colleagues developed the Model Authoring Toolkit to help practitioners consider diverse stakeholder values and deliberate on system design trade-offs (Shen et al., 2022). Their qualitative and survey study of Wikipedia communities found that framing ethical reflection and documentation as a participatory process led to more informed decisions about AI system design and deployment.
Challenges. The assumption that documentation artifacts and processes can promote ethical reasoning among practitioners hinges on the belief that practitioners are sufficiently aware of AI’s ethical risks to users, non-users, and society that they can recognize them when prompted. However, without proper training in responsible AI practices and exposure to groups potentially harmed by AI systems, practitioners may not connect documentation with ethical risks (Heger et al., 2022). Consequently, their consideration of harms may only partially encompass the scope of risks, especially those requiring non-technical mitigations. For example, practitioners often misunderstand how bias can manifest in their work, leading them to incorrectly state in their documentation that bias concerns are not relevant (Hind et al., 2019). While it may not be necessary for ethicists to be embedded in the work process for practitioners to make responsible choices, as some have suggested (McLennan et al., 2022), findings point to the need for more support and training for practitioners in how to identify, document, and effectively communicate potential risks (Heger et al., 2022; Madaio et al., 2024).
Furthermore, practitioners and organizations may resist documenting ethical impacts they have identifed. Publicly available documentation artifacts, such as Hugging Face model cards and GitHub repository README files, often show little consideration of ethical concerns (Pepe et al., 2024; Liang et al., 2024). While this omission might occur because practitioners are unaware of potential harms (Heger et al., 2022), their absence could also result from practitioners being hesitant to record ethical concerns they did recognize. Sometimes practitioners resist addressing ethical impacts because they feel unqualified to speculate on numerous potential use cases and their impacts (Crisan et al., 2022; Heger et al., 2022; Hind et al., 2019; Nunes et al., 2022). Others worry that detailing ethical concerns might give downstream stakeholders a false sense of security (Crisan et al., 2022). This hesitation could also stem from a desire to avoid personal responsibility for their actions (Nunes et al., 2022). Alternatively, practitioners could plausibly have concerns that documenting potential ethical harms could create a legal or public relations risk if outside stakeholders gained access to documentation.
Even if documentation effectively engages practitioners in ethical reflection, they may have little influence over their organization’s development goals, deliverables, and timelines. As a result, they might be unable to make the changes to data, models, or systems that they have identified as useful or necessary to address ethical risks through the documentation process. One ethnographic study found that business demands, not the beliefs of data subjects or practitioners, largely determined the organization’s documentation approaches (Miceli et al., 2021). For example, although practitioners who reflected on data labeling practices recognized that socially constructed identities are complex, the organization nevertheless chose to represent identity in a reductive way, such as by defining race according to discrete, mutually exclusive categories. The authors concluded that explicit and implicit power structures among internal organizational stakeholders significantly affect practitioners’ documentation practices and ability to shape outcomes. Documentation approaches must be responsive to these constraints. Otherwise, documentation about ethical considerations may not promote meaningful action or may be incomplete or misleading.
On the whole, evidence on the extent to which AI documentation frameworks can truly enhance ethical decision-making is mixed. Practitioners are capable of ethical deliberation about AI, but do not always thoroughly explore ethical issues or record the results in documentation artifacts. Even if practitioners engage in ethical deliberation when creating and using documentation, organizational environments can enable or constrain individual practitioners’ normative choices.
4.2.4. Potential impact #4: Documentation could catalyze overall improvements in governance and development.
Researchers have suggested that documentation can significantly influence governance and development practices broadly. In practical settings, AI documentation doesn’t function independently; documentation processes typically require practitioners to engage in other activities that can improve AI outcomes. Thus, the documentation process can catalyze behaviors that enhance the quality of AI system development and governance beyond documentation processes and artifacts.
Hypotheses. Framework authors posit that documentation could improve the rigor with which practitioners develop AI systems (Holland et al., 2018). By requiring practitioners to justify their development choices, documentation may lead to more careful decision-making (Bender and Friedman, 2018; Chmielinski et al., 2024; Holland et al., 2018). Also, clear and comprehensive documentation supports reproducibility, aiding other practitioners in retraining and validating systems consistently (Adkins et al., 2022; Baracaldo et al., 2022; Hind et al., 2019). Since AI practitioners are often driven by a commitment to scientific rigor (Winecoff and Watkins, 2022), documentation that emphasizes this aspect can serve as a ”value lever (Shilton, 2013),” further encouraging engagement with the documentation process .
Documentation might also improve development and governance efficiency. In large organizations, datasets, models, and systems developed for one purpose may also be useful to other teams with similar goals. Documentation can make system components more discoverable, helping practitioners avoid duplicating efforts (Heger et al., 2022). Documentation can quickly provide information on the limitations of using a dataset, model, or system (Rostamzadeh et al., 2022), which could allow practitioners to allocate more time to identifying and addressing any additional risks that may arise when integrating components into new systems. It can also include details relevant to company or legal policies that apply to system components, preventing practitioners from wasting time on products and features that their organizations ultimately won’t approve for deployment (Roman et al., 2023). Practitioners also point out that by documenting unsuccessful and successful approaches, organizations can prevent repeated mistakes and identify practices that would be beneficial for the organization to disseminate broadly (Chang and Custis, 2022; Miceli et al., 2021).
Documentation could also facilitate proactive risk mitigation, which is often more efficient than reactive approaches (Mitchell et al., 2019). When issues arise in already-deployed systems, organizations typically need to fix the problem without disrupting service quality or availability. They might need to roll back to a previous version of the system that is less performant or apply quick patches that may not fully resolve the issue. Documentation that helps identify potential problems before deployment can lead to more comprehensive solutions than addressing issues post hoc once they are discovered in production systems (Holland et al., 2018). And helping teams discover and use approved AI components that have already undergone relevant risk management processes rather than creating new ones can help ensure risks are not re-introduced and overlooked.
Furthermore, documentation could preserve institutional knowledge. In any organization, particularly complex ones, consistently institutionalizing values and best practices can be challenging. Documentation could serve as a means of conveying technical information about systems, and the organization’s policies and values that influence development and use. For instance, documentation can aid in onboarding new employees by providing information about the systems they’ll work on and communicating the organization’s approach to development and governance (Adkins et al., 2022; Chang and Custis, 2022; Heger et al., 2022).
Lastly, documentation could facilitate both internal and external audits (Brajovic et al., 2023; Crisan et al., 2022; Hupont and Gomez, 2022; Miceli et al., 2021). Audits may focus on the components of AI systems, the overall application, or the processes used in development (Mökander et al., 2023). Documentation could help organizations demonstrate that their claims—such as providing equitable system performance across different demographic groups—are backed by evidence. In these cases, documentation could guide auditors in selecting appropriate methods based on the auditor’s objectives and the available data. When audits examine the robustness of an organization’s governance, documentation becomes even more critical (Clavell, 2024). Without comprehensive documentation, organizations may struggle to demonstrate the integrity and consistency of their processes.
Challenges. One barrier to documentation facilitating best practices in development and governance is that the documentation process must be interconnected with other organizational processes to serve as an effective forcing function. In cases where documentation is ad-hoc and relies on tools that are not well-integrated with practitioners’ workflows (Bhat et al., 2023; Heger et al., 2022), connecting documentation processes and artifacts to other aspects of the development and governance process becomes even more challenging. Without a solid documentation infrastructure that aligns with other organizational functions, documentation may fail to effectively enhance development and risk management.
Another challenge is that while high-quality documentation can improve governance and development, poor-quality documentation can have negative impacts. Documentation should prevent redundant efforts, maximize time spent managing unique risks, and support proactive system development. However, these benefits depend on the quality of the documentation. Inadequate or incorrect documentation can lead to wasted time, overlooked risks, and unexpected issues that may require correction later in development. This could result in harms remaining unaddressed or inefficiencies that undermine the overall governance process.
4.3. Design considerations
How organizations choose to document AI systems reflects both their normative values and practical constraints. Design and implementation choices are also influenced by established and emerging external requirements, such as regulations that companies must adhere to. As organizations (and regulators) aim to maximize the benefits of documentation across various types of AI systems and components, they must often navigate tensions in their approach. Our analysis identified common design tensions or considerations organizations must navigate. These include: balancing customization and standardization, tailoring documentation to specific audiences, determining the appropriate level of detail, deciding on the degree of automation in the documentation process, and incorporating interactivity into documentation artifacts. To an extent, these design considerations are interrelated: the level of detail documentation artifacts contain can also be a form of audience tailoring, for example. Yet each of these design tensions has partially distinct motivations and implications, so we address them separately.
4.3.1. Design consideration #1: Determining the degree of customization versus standardization in documentation artifacts.
Organizations often create datasets, models, and systems for various purposes, employing a wide range of techniques and components. This diversity leads to a variety of risks, which customization is well-suited to address. At the same time, standardized documentation helps establish consistent practices and institutionalize norms.
One advantage of customization is that it can address the specific risks and capabilities of an organization or AI system. For instance, annotating the origin of data collected in different healthcare settings can be crucial, as different settings serve distinct patient populations and employ varied healthcare practitioners. This information can help practitioners identify gaps in the dataset or regions where the dataset may be less applicable (Rostamzadeh et al., 2022). Conversely, annotating the origin of code snippets from software engineers working in different contexts may be less important since code typically functions similarly across different environments.
Echoing calls to ground approaches to AI harms within the context of deployment (Hutchinson et al., 2022; Narayanan and Kapoor, 2024; Nicholas, 2024), several frameworks recommend collectively documenting the system components that pertain to a given use case rather than separately documenting datasets, models, or systems independent of their intended use case (Chmielinski et al., 2020; Hupont and Gomez, 2022; Mohammad, 2022; Rostamzadeh et al., 2022; Srinivasan et al., 2021). For example, some researchers have proposed specific documentation methods for affective computing (Hupont and Gomez, 2022), since these methods pose unique risks related to psychological manipulation and surveillance and so may merit a tailored approach. Moreover, because affective computing applications are likely to be classified as high-risk by the EU AI Act, applications using affective computing are also more likely to be subject to specific documentation requirements (Hupont and Gomez, 2022).
Customization can also be valuable at the model level, as different models are designed to perform unique functions and thus pose unique risks. In one study, practitioners argued that the information within a model card should be rearranged to ensure that documentation consumers clearly understand the specific model’s purpose and limitations (Crisan et al., 2022). Another study found that practitioners added technical details to standardized documentation, even when instructed not to, suggesting that they believe standardized formats need flexibility to include model-specific details they find relevant (Bhat et al., 2023).
At the organizational level, customization might also be necessary. For example, a healthcare startup may require a different documentation approach than a large, well-established financial institution. In some cases, it may be appropriate to customize documentation at multiple levels to respond to both the organization’s nature and the type of system it is developing (Chmielinski et al., 2020; Gebru et al., 2021; Holland et al., 2018; Richards et al., 2020; Roman et al., 2023; Stoyanovich and Howe, 2019).
While customization has many benefits, it also presents challenges. Standard formats facilitate the development of tools that allow practitioners to search for system components that meet particular specifications, improving discoverability and efficiency within an organization’s ecosystem (Heger et al., 2022). Since such tools might enable practitioners to more easily find appropriate components, it might also reduce the likelihood that they employ inappropriate ones. This enhanced discoverability is often cited as one of the most significant benefits of documentation by practitioners (Heger et al., 2022; Reid and Williams, 2023).
Standardized documentation could also facilitate practitioners comparing candidate datasets, models, and systems, and for comparing AI and non-AI approaches (Chmielinski et al., 2024; Gilbert et al., 2023; Hind et al., 2019; Mitchell et al., 2019; Reid and Williams, 2023; Shen et al., 2022; Stoyanovich and Howe, 2019; Zheng et al., 2022). When documentation artifacts vary significantly in the type of information they contain or how this information is presented, comparisons become more challenging. Given the time constraints and the currently limited incentives for creating (Chang and Custis, 2022; Heger et al., 2022; Miceli et al., 2021) and using (Crisan et al., 2022) documentation, maintaining the core utility of documentation through standardization could support broader adoption within organizations.
A more diffuse but critical benefit of standardization is its role in helping the AI community converge on norms of practice and communication (Arnold et al., 2019; McMillan-Major et al., 2021, 2024). Practitioners often struggle with how to complete documentation or what level of detail to include (Chang and Custis, 2022; Pushkarna et al., 2022; Heger et al., 2022; Miceli et al., 2021), leading them to rely heavily on existing examples, which may not always be appropriate. In one study, practitioners frequently copied documentation from previous projects into their current work, sometimes including irrelevant information or omitting crucial details (Pushkarna et al., 2022). Standardization can mitigate these issues by providing clear expectations for what information should be included and what users should expect to find in documentation artifacts, thus encouraging more consistent practices (McMillan-Major et al., 2021; Pushkarna et al., 2022; Roman et al., 2023).
Standardization also might facilitate structured risk management. If documentation producers are instructed to select certain details from structured categories, or certain fields are constrained to particular structured formats, these fields can be used to trigger certain governance actions, such as scheduling a review or requiring that a certain mitigation be applied prior to proceeding. On the other hand, such process automation can mean that concepts and risks are oversimplified, and that processes or mitigations are recommended in cases where they may not be appropriate, while leaving other relevant risks to be overlooked.
The choice between customization and standardization is not always a strict binary but a spectrum, allowing for a combination of both. For example, organizations can create standardized templates that are adapted to different contexts, such as healthcare models versus those for software engineers, ensuring consistency while meeting specific needs. However, organizations must thoughtfully decide when to standardize and when to customize.
4.3.2. Design consideration #2: Determining how much to tailor documentation artifacts to specific audiences
When designing documentation strategies, organizations must consider the diverse needs of multiple stakeholder groups. Different stakeholders often require distinct forms of documentation to be effective. To meet these varied needs, organizations might choose to produce tailored documentation for each stakeholder group. Alternatively, organizations might opt for a single, general-purpose artifact that all stakeholders can use, though this might not be adequately tailored to a specific purpose.
A single-format artifact may not adequately serve any particular stakeholder group. For instance, documentation designed for data scientists might not be accessible or useful to UX professionals or other non-technical stakeholders (Liao et al., 2023). Similarly, documentation aimed at public transparency might lack the depth needed by internal decision-makers handling risk management. This highlights a key trade-off: while a single-format document can reduce the complexity of documentation management, it risks failing to meet the specialized needs of different groups.
Stakeholder-specific documentation, meanwhile, can address the unique needs of different groups of practitioners. These groups, including users, policymakers, data scientists, engineers, lawyers, and others (Micheli et al., 2023), often have distinct goals and use different terminologies, which can influence the type of documentation produced and the most helpful forms for each group. Documentation tailored to its specific audience is more likely to provide stakeholders with the necessary information for accomplishing their goals. Without a clear target audience, documentation producers might only address what is relevant to their own team (Heger et al., 2022), potentially overlooking critical information needed by other teams (Hind et al., 2019) or presenting it in an unusable format (Liao et al., 2023).
Tailoring documentation is particularly important when conveying technical information to non-technical stakeholders, both within and outside the organization. While many researchers propose frameworks that aim to make documentation comprehensible to both technical and non-technical practitioners (Holland et al., 2018; McMillan-Major et al., 2021; Mohammad, 2022; Pushkarna et al., 2022; Richards et al., 2020; Roman et al., 2023; Sokol and Flach, 2020; Tagliabue et al., 2021), empirical studies indicate that non-technical audiences often struggle to understand even simplified technical details (Crisan et al., 2022; Liao et al., 2023; Miceli et al., 2021; Shen et al., 2022). For example, in a study where non-technical practitioners used documentation to assess model risks, a significant proportion failed to correctly identify the model’s basic purpose (Crisan et al., 2022). Using documentation, non-technical practitioners may struggle to grasp technical information, such as accuracy metrics (Shen et al., 2022). This suggests that unless technical information is presented in simpler terms, non-technical stakeholders may still face challenges in interpreting technical information necessary for risk assessment when it is not specifically tailored to their needs.
Another advantage of stakeholder-specific documentation is its ability to leverage the expertise of different practitioners more effectively. Tailoring documentation allows stakeholders to focus on the most relevant information for their roles, thereby enhancing the organization’s ability to identify and manage risks. For instance, in developing a documentation framework for healthcare data, machine learning experts prioritized the dataset’s composition, while healthcare experts focused on the details of medical diagnoses and data collection processes, as these were crucial for interpreting the data’s relevance and limitations (Rostamzadeh et al., 2022). By emphasizing information pertinent to their expertise, stakeholders can conduct more thorough analyses and potentially uncover issues that others might miss (Pushkarna et al., 2022; Rostamzadeh et al., 2022).
Tailored documentation can also be more actionable for practitioners. One common issue with general-purpose documentation is that it may not clearly guide stakeholders on how to use the information provided (Adkins et al., 2022; Crisan et al., 2022; Stoyanovich and Howe, 2019). This issue is particularly significant when practitioners fail to connect documentation with the ethical impacts of their work (Heger et al., 2022), indicating a need for prescriptive guidance to manage AI risks effectively. The actions practitioners take in response to documentation often depend on their specific roles for two reasons. First, when documentation is aligned with the purpose of their role, practitioners are better able to understand the necessary actions to manage risk. For example, information about how well a system meets user needs may be more actionable for a UX professional than a machine learning engineer. Second, practitioners typically have the authority to act in specific ways within organizations. For instance, compliance teams cannot implement changes to the production codebase, and data scientists are not responsible for conducting legal reviews. Tailored documentation that aligns with each group’s roles and responsibilities can guide them toward actions they are empowered to take.
Whereas multiple forms of documentation can better meet diverse stakeholder needs, single-format documentation can help reduce confusion and avoid the fragmentation that occurs when information is scattered across different formats, such as README files, wikis, and slide decks (Heger et al., 2022). Even with centralized repositories that link different versions, managing multiple formats can still lead to confusion about the existence and authority of these artifacts and make it difficult for practitioners to access the information they need and develop a cohesive understanding.
Furthermore, single-format documentation may be easier to produce and maintain. If practitioners are required to create multiple forms of documentation to meet the needs of different stakeholders, organizations must ensure sufficient time is allocated for this process. Without adequate time, practitioners may rush, leading to low-quality documentation. In cases where time constraints are a concern, it may be more effective to focus on producing a single, high-quality document that is regularly updated. At a minimum, all documentation should clearly indicate its last update to help users gauge its relevance and accuracy.
4.3.3. Design trade-off #3: Determining the level of interactivity documentation should support
Producing multiple forms of documentation is one way to support a variety of goals, stakeholders, and systems and goals; however, these multiple forms of documentation create challenges for organizations to manage. Interactive documentation offers an alternative solution by enabling stakeholders to access the specific information they need while maintaining overall usability and comprehension.
Interactive documentation may be able to effectively serve multiple stakeholders with a single artifact. For example, interactive system diagrams might allow practitioners to click on specific components to access technical details (Procope et al., 2022), while expandable sections can provide additional context as necessary (Crisan et al., 2022).
Empirical research supports the idea that interactive documentation can significantly improve understanding, particularly among non-technical audiences (Crisan et al., 2022; Liao et al., 2023; Moore et al., 2023). Tools like Hugging Face’s model inference API444https://huggingface.co/docs/api-inference/ or OpenAI’s developer playground555https://platform.openai.com/playground allow practitioners to engage with models and systems directly, fostering an intuitive understanding of how inputs and outputs are connected. This hands-on interaction is crucial for designing products that leverage the system’s capabilities and for assessing potential risks (Liao et al., 2023). Interactive access is particularly valuable when exploring pre-trained, general-purpose models, whose functionality can vary significantly depending on the implementation (Liao et al., 2023; Moore et al., 2023; Crisan et al., 2022).
However, interactive documentation also presents challenges. Without good information architecture, it can overwhelm users, leading to cognitive overload rather than enhancing understanding (Crisan et al., 2022). Additionally, interactive documentation requires additional engineering effort to design and maintain, which may be resource-intensive for organizations. In contrast, static documentation offers consistency and simplicity. It presents information in a uniform format, fostering a common understanding that can enhance communication, collaboration, and decision-making across teams.
4.3.4. Design consideration #4: Determining the level of detail documentation artifacts should contain.
Organizations must determine what level of detail to include in documentation artifacts. Whereas exhaustive documentation might ensure practitioners have access to all the necessary information, concise documentation may improve ease of production and use.
Concise documentation can help stakeholders focus on the details that are most relevant to their needs, reducing the likelihood of information overload. For documentation aimed at downstream consumers, including internal risk management professionals, it is essential to provide enough detail to support decision-making without overwhelming them. Information overload can lead to selective attention, where decision-makers may focus on certain details while neglecting others, potentially degrading the quality of their decisions (Phillips-Wren and Adya, 2020), particularly under time constraints (Hahn et al., 1992). For example, faced with extensive and complex documentation on AI systems, consumers might miss critical information or become frustrated and abandon the effort. As one participant in a research study noted, ”I would lose patience after 30-40 seconds if I have to put a lot of effort into finding what I’m looking for” (Crisan et al., 2022). Concise documentation that highlights the most important information for a given stakeholder group can help internal decision-makers manage risk more effectively by ensuring they focus on critical details.
Moreover, concise documentation is often easier to produce and maintain, which can enhance accuracy and encourage more frequent use. Exhaustive documentation may be burdensome to create and keep updated, particularly if the process is manual. For instance, in the initial case studies of an extensive data documentation framework, practitioners reported that completing the documentation took two to three hours, excluding the time needed to gather the required information (Bender and Friedman, 2018). While a few hours might be manageable for a single project, this time commitment becomes overwhelming when scaled across an entire organization, potentially involving hundreds or thousands of datasets, models, or systems. In such scenarios, spreading effort thinly across all systems might detract from the focus on high-risk or high-impact components, undermining critical risk management efforts.
However, while concise documentation has its advantages, it may omit details that, although not relevant to most uses, are crucial for specific applications. Exhaustive documentation provides a more comprehensive overview that can accommodate a broader range of applications. For example, consider a dataset containing patient information and healthcare outcomes over a year. If a model uses this dataset to predict patients’ health outcomes over time, the model’s accuracy depends on the validity and reliability of the data’s timestamps. Documenting how these timestamps were applied is essential for practitioners intending to use the dataset for this purpose. On the other hand, if the dataset is used for clustering patients into broad categories, temporal information becomes less critical. Therefore, the project’s purpose and the specific stakeholders involved should shape the level of detail included in the documentation.
4.3.5. Design consideration #5: Determining the amount of automation in the documentation process.
Documentation encompasses two main types of information: data that can be extracted automatically from the system and information that requires human input. Information that could potentially be derived directly from source code or generated through scripts interacting with system components includes elements such as data quantity, distribution, statistical properties, model types and parameters, system flow, and software library usage. In contrast, information related to human decision-making processes and organizational context requires manual input. This includes the motivations for developing specific datasets or models, criteria for selecting data sources, reasoning behind chosen methodologies, outcomes of compliance reviews, contact information for responsible parties, decommissioning procedures, and ethical considerations addressed during development.
One of the purposes of both producing and using documentation within an organization is to prompt developers to consider the ethical implications of their systems. However, the degree of automation in the documentation process can significantly influence how thoroughly practitioners engage with these considerations. Fully automated documentation requires minimal direct involvement, which may not encourage the level of critical reflection necessary to address complex ethical issues. As a result, some experts advocate for completing documentation manually, even when automation is feasible, to ensure that practitioners engage deeply with the material (Gebru et al., 2021; McMillan-Major et al., 2024).
However, automated documentation offers significant benefits in terms of efficiency and accuracy. Manually creating documentation is time-consuming and prone to errors, particularly if it is not well integrated with practitioners’ existing tools and workflows (Bhat et al., 2023). Practitioners have noted that automation can reduce the time required to complete documentation and increase the likelihood that it is kept up-to-date (Heger et al., 2022). Automation also reduces the burden on practitioners to recall critical information after the fact (Hind et al., 2019), which can be particularly valuable in complex, fast-paced environments. Organizations should also be aware that practitioners facing multiple competing priorities may circumvent manual processes by implementing their own forms of unaccountable automation. Well-designed automation of some portions of documentation could reduce this risk.
4.4. Recommendations
Given that organizations have unique requirements and operate in vastly different contexts, there is no one-size-fits-all approach to documentation or a universal set of information that applies to all datasets, models, or systems. However, our analysis of proposed frameworks and empirical findings suggests several considerations that organizations should take into account when designing and implementing their documentation strategies. Each of these recommendations could be further strengthened by novel empirical research about how organizations can maximize the benefits of implementing the recommendations in applied contexts.
Identify critical stakeholders and their documentation needs. When developing AI documentation strategies, organizations should begin by identifying the stakeholders who will frequently and significantly interact with both the documentation processes and artifacts. It’s important to consider not only the needs of those who will consume the documentation but also the expertise and capacity of those responsible for producing it. If the documentation is designed to be highly informative for downstream users but is impractical for upstream practitioners to produce, the approach is likely to fail in practice. In such cases, organizations may need to expand their documentation teams to ensure they can meet the needs of consumers effectively. Also, organizations should take into account the technical and non-technical backgrounds of their stakeholders to strike a balance between providing the necessary level of detail and maintaining accessibility.
Evaluate operational resources and constraints and incentivize the production of robust documentation within that context. A realistic assessment of available resources—such as time, personnel, and expertise—is crucial for supporting effective documentation. Overly ambitious documentation efforts without adequate resources can lead to failures, such as incomplete or inaccurate documentation. It is often better to produce concise, high-quality documentation than to attempt a comprehensive approach that falls short. Organizations should ensure that practitioners understand their documentation responsibilities and that incentives are in place to promote thorough and accurate documentation practices.
Leverage empirical evidence and feedback for continuous improvement. Organizations should evaluate potential documentation approaches via pilot projects and periodically collect empirical data on the efficacy of their documentation practices in meeting the needs of relevant stakeholders. Organizations should plan to assess the usability of documentation frameworks and their effectiveness in achieving the organization’s goals (Berman et al., 2024), and adapt their documentation strategies by conducting continuous assessments as their needs or technology evolves. Given the limited empirical data on the effectiveness of documentation approaches, organizations can also contribute to the broader AI community by publicly sharing their findings on what works best in different contexts. Organizations should recognize that empirical evidence and key performance indicators can be qualitative and quantitative. Whereas quantitative evidence can help organizations assess general trends at scale, qualitative evidence can help them understand unique issues in greater depth.
Use standardization as the default approach, and intentionally deviate from it when necessary. Standardization offers many advantages. It allows practitioners to compare documentation artifacts, facilitates search and discoverability of information, establishes consistent expectations for documentation producers and consumers, and can help institutionalize norms of practice. For these reasons, some level of standardization in documentation processes and artifacts is likely to be helpful for most organizations for most AI systems.
To balance standardization and customization, organizations might develop customizable documentation templates or interchangeable modules (Bhat et al., 2023; Heger et al., 2022; Hind et al., 2019; Holland et al., 2018; McMillan-Major et al., 2021; Pushkarna et al., 2022; Richards et al., 2020). Templates could, for example, allow for customization of subsections while maintaining consistent content and structure within those subsections. For instance, companies may include a subsection on the labeling process for AI systems based on supervised learning but exclude it for unsupervised AI systems since they do not require labels. Where companies choose to embrace a greater degree of customization, they should ensure that the benefits of this customization outweigh the benefits of maintaining standardization.
Leverage automation for information extraction and incorporate manual processes for scoped critical synthesis. Some helpful information for documentation, such as dataset composition, statistical properties, model training methods, parameters, and software library or third-party API usage, can be extracted from AI systems using automated procedures, contributing to consistent, high-fidelity documentation of these system elements. While automated documentation can provide reliable and up-to-date information, it has limitations. The aspects of systems that are easiest to document automatically are not always the most critical, and errors in automated documentation can have a greater negative impact than manual errors, as they can persist through each update. Therefore, organizations should be selective about the information included in documentation artifacts, even if it can be easily generated automatically.
Consider interactive documentation interfaces for general-purpose models in particular. Documenting the characteristics of general-purpose models and their potential risks is particularly challenging due to their wide range of possible applications. When downstream practitioners cannot directly communicate with upstream model developers, interpreting the general documentation in the context of specific applications can be difficult. Several studies have shown that interfaces allowing practitioners to observe system outputs in response to given inputs can enhance their understanding of model capabilities and risks (Liao et al., 2023; Crisan et al., 2022; Moore et al., 2023). Consequently, interactive interfaces can serve as a valuable complement to traditional, static documentation.
Focus initially on core information needs for documentation and evaluate potential gaps over time. When designing and implementing new documentation strategies, organizations should start by identifying the essential pieces of information that will meet the needs of key stakeholders. In other words, they should start by designing a concise version of documentation. They should plan to consult with stakeholders who use this documentation to determine whether any additional information is necessary to perform their essential duties and whether any of the initial information provided in documentation artifacts is unnecessary. They should also incorporate learnings from empirical research to augment documentation as needed. Organizations can use this agile approach to manage the information in documentation artifacts, ensuring that the time spent producing and using documentation is maximally efficient and effective at serving stakeholder requirements.
Institutionalize responsible AI and risk management practices beyond documentation. Ethical deliberation is a skill that practitioners can develop over time with intentional practice (Vallor et al., 2018), but organizations need to implement additional interventions beyond documentation to help cultivate this skill. Even when practitioners engage in documentation processes, they may overlook critical risks or fail to recommend appropriate mitigations without a deeper understanding of how AI can harm diverse individuals and communities (Madaio et al., 2024; Chang and Custis, 2022). Therefore, organizations should strive to embed responsible AI and risk management efforts throughout their AI development practices. When responsible AI is integrated into the organization’s culture and broader practices, documentation processes that prompt practitioners to reflect on risks are more likely to be informed by a comprehensive understanding necessary for effective risk management. Conversely, relying solely on documentation to encourage critical reflection is unlikely to be sufficient and may even lead to negative outcomes. Documentation of potential ethical harms will likely be more effective when it focuses on concrete system properties, such as data containing personally identifiable information or existing societal biases, rather than on speculative concerns, which can be subjective and heavily influenced by the expertise and perspectives of the documentation producers.
4.5. Conclusion
This paper underscores the critical role of documentation in enhancing AI governance within organizations, emphasizing that effective documentation practices are essential for managing AI risks and fostering responsible system development. To support robust governance, organizations must tailor their documentation processes and artifacts to meet the specific needs and constraints of their stakeholders. Establishing clear success criteria—such as accuracy, comprehensiveness, and usability—and regularly assessing progress against these goals is crucial for maintaining the effectiveness of documentation strategies.
While our focus has been on documentation for internal governance, these processes and artifacts can also form the foundation for transparency efforts directed at external stakeholders. By cultivating strong internal documentation practices and continually evaluating their success, organizations can build the necessary infrastructure to create detailed, actionable records of system development and risk management. These records are necessary for communicating effectively with external audiences, such as regulators and the public.
However, documentation designed for internal use may not seamlessly translate to external contexts. The differences in expertise, needs, and objectives between internal and external stakeholders require careful consideration. Organizations must adapt their documentation processes and artifacts to bridge these gaps, ensuring that they are both accessible and informative to all relevant parties. By doing so, organizations can enhance their governance efforts and contribute to a more transparent and accountable AI ecosystem.
Acknowledgements.
We would like to thank members of the AI Governance Lab team at CDT and Emily McReynolds for their valuable feedback on our research design, analysis, and synthesis. We also extend our thanks to Samir Jain and Drew Courtney for their insightful comments on our manuscript. Finally, we are grateful to the participants of the ”Improving Documentation for AI Governance” workshop, held at CDT in June 2024.References
- (1)
- Adkins et al. (2022) David Adkins, Bilal Alsallakh, Adeel Cheema, Narine Kokhlikyan, Emily McReynolds, Pushkar Mishra, Chavez Procope, Jeremy Sawruk, Erin Wang, and Polina Zvyagina. 2022. Prescriptive and Descriptive Approaches to Machine-Learning Transparency. In CHI Conference on Human Factors in Computing Systems Extended Abstracts. 1–9. https://doi.org/10.1145/3491101.3519724
- Ahlawat et al. (2024) Archana Ahlawat, Amy Winecoff, and Jonathan Mayer. 2024. Minimum Viable Ethics: From Institutionalizing Industry AI Governance to Product Impact. arXiv (September 11 2024). https://www.arxiv.org/abs/2409.06926
- Ali et al. (2023) Sanna J Ali, Angèle Christin, Andrew Smart, and Riitta Katila. 2023. Walking the Walk of AI Ethics: Organizational Challenges and the Individualization of Risk among Ethics Entrepreneurs. In 2023 ACM Conference on Fairness, Accountability, and Transparency. ACM, Chicago IL USA, 217–226. https://doi.org/10.1145/3593013.3593990
- Arnold et al. (2019) Matthew Arnold, Rachel K. E. Bellamy, Michael Hind, Stephanie Houde, Sameep Mehta, Aleksandra Mojsilovic, Ravi Nair, et al. 2019. FactSheets: Increasing Trust in AI Services through Supplier’s Declarations of Conformity. arXiv (February 7 2019). http://arxiv.org/abs/1808.07261
- Baracaldo et al. (2022) Nathalie Baracaldo, Ali Anwar, Mark Purcell, Ambrish Rawat, Mathieu Sinn, Bashar Altakrouri, Dian Balta, et al. 2022. Towards an Accountable and Reproducible Federated Learning: A FactSheets Approach. arXiv (February 24 2022). http://arxiv.org/abs/2202.12443
- Belz et al. (2023) Anya Belz, Craig Thomson, Ehud Reiter, Gavin Abercrombie, Jose M. Alonso-Moral, Mohammad Arvan, Anouck Braggaar, et al. 2023. Missing Information, Unresponsive Authors, Experimental Flaws: The Impossibility of Assessing the Reproducibility of Previous Human Evaluations in NLP. arXiv (August 7 2023). http://arxiv.org/abs/2305.01633
- Bender and Friedman (2018) Emily M. Bender and Batya Friedman. 2018. Data Statements for Natural Language Processing: Toward Mitigating System Bias and Enabling Better Science. Transactions of the Association for Computational Linguistics 6 (December 2018), 587–604. https://doi.org/10.1162/tacl_a_00041
- Berman et al. (2024) Glen Berman, Nitesh Goyal, and Michael Madaio. 2024. A Scoping Study of Evaluation Practices for Responsible AI Tools: Steps Towards Effectiveness Evaluations. In Proceedings of the CHI Conference on Human Factors in Computing Systems. ACM, Honolulu HI USA, 1–24. https://doi.org/10.1145/3613904.3642398
- Bhat et al. (2023) Avinash Bhat, Austin Coursey, Grace Hu, Sixian Li, Nadia Nahar, Shurui Zhou, Christian Kästner, and Jin L. C. Guo. 2023. Aspirations and Practice of Model Documentation: Moving the Needle with Nudging and Traceability. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–17. https://doi.org/10.1145/3544548.3581518
- Blasch et al. (2021) Erik Blasch, James Sung, and Tao Nguyen. 2021. Multisource AI scorecard table for system evaluation. In AAAI FSS20: Artificial Intelligence in Government and Public Sector.
- Boyd (2021) Karen L. Boyd. 2021. Datasheets for Datasets Help ML Engineers Notice and Understand Ethical Issues in Training Data. Proceedings of the ACM on Human-Computer Interaction 5, CSCW2 (October 13 2021), 1–27. https://doi.org/10.1145/3479582
- Brajovic et al. (2023) Danilo Brajovic, Niclas Renner, Vincent Philipp Goebels, Philipp Wagner, Benjamin Fresz, Martin Biller, Mara Klaeb, et al. 2023. Model Reporting for Certifiable AI: A Proposal from Merging EU Regulation into AI Development. arXiv (July 21 2023). http://arxiv.org/abs/2307.11525
- Chang and Custis (2022) Jiyoo Chang and Christine Custis. 2022. Understanding Implementation Challenges in Machine Learning Documentation. In Equity and Access in Algorithms, Mechanisms, and Optimization. ACM, Arlington VA USA, 1–8. https://doi.org/10.1145/3551624.3555301
- Charmaz (2014) Kathy Charmaz. 2014. Constructing grounded theory, 2nd ed. Sage.
- Chmielinski et al. (2024) Kasia Chmielinski, Sarah Newman, Chris N Kranzinger, Michael Hind, Jennifer Wortman Vaughan, Margaret Mitchell, Julia Stoyanovich, et al. 2024. The CLeAR Documentation Framework for AI Transparency Recommendations for Practitioners and Context for Policymakers. The Shorenstein Center on Media, Politics and Public Policy (May 2024). https://shorensteincenter.org/clear-documentation-framework-ai-transparency-recommendations-practitioners-context-policymakers/
- Chmielinski et al. (2020) Kasia S Chmielinski, Sarah Newman, Matt Taylor, Josh Joseph, Kemi Thomas, Jessica Yurkofsky, and Yue Chelsea Qiu. 2020. The Dataset Nutrition Label (2nd Gen): Leveraging Context to Mitigate Harms in Artificial Intelligence.
- Clavell (2024) Gemma Galdon Clavell. 2024. Checklist for AI Auditing. https://www.edpb.europa.eu/system/files/2024-06/ai-auditing_checklist-for-ai-auditing-scores_edpb-spe-programme_en.pdf
- Crisan et al. (2022) Anamaria Crisan, Margaret Drouhard, Jesse Vig, and Nazneen Rajani. 2022. Interactive Model Cards: A Human-Centered Approach to Model Documentation. In 2022 ACM Conference on Fairness, Accountability, and Transparency. 427–439. https://doi.org/10.1145/3531146.3533108
- Díaz et al. (2023) Mark Díaz, Sunipa Dev, Emily Reif, Emily Denton, and Vinodkumar Prabhakaran. 2023. SoUnD Framework: Analyzing (So)Cial Representation in (Un)Structured (D)Ata. arXiv (December 1 2023). http://arxiv.org/abs/2311.17259
- Friedman et al. (2006) Batya Friedman, Peter H. Jr. Kahn, and Alan Borning. 2006. Value Sensitive Design and Information Systems. In Human-Computer Interaction and Management Information Systems: Foundations Advances in Management Information Systems. M.E. Sharpe, Armonk, NY, 348–372. http://ndl.ethernet.edu.et/bitstream/123456789/18272/1/87..Neelke%20Doorn.pdf#page=67
- Gebru et al. (2021) Timnit Gebru, Jamie Morgenstern, Briana Vecchione, Jennifer Wortman Vaughan, Hanna Wallach, Hal Daumé III, and Kate Crawford. 2021. Datasheets for Datasets. arXiv (December 1 2021). http://arxiv.org/abs/1803.09010
- Geiger et al. (2020) R Stuart Geiger, Kevin Yu, Yanlai Yang, Mindy Dai, Jie Qiu, Rebekah Tang, and Jenny Huang. 2020. Garbage in, garbage out? Do machine learning application papers in social computing report where human-labeled training data comes from?. In Proceedings of the 2020 conference on fairness, accountability, and transparency. 325–336.
- Gilbert et al. (2023) Thomas Krendl Gilbert, Nathan Lambert, Sarah Dean, Tom Zick, and Aaron Snoswell. 2023. Reward Reports for Reinforcement Learning. arXiv (March 19 2023). http://arxiv.org/abs/2204.10817
- Green et al. (2021) Nekesha Green, Chavez Procope, Adeel Cheema, and Adekunle Adediji. 2021. System Cards, a New Resource for Understanding How AI Systems Work. https://ai.facebook.com/blog/system-cards-a-new-resource-for-understanding-how-ai-systems-work/
- Hahn et al. (1992) Minhi Hahn, Robert Lawson, and Young Gyu Lee. 1992. The Effects of Time Pressure and Information Load on Decision Quality. Psychology & Marketing 9, 5 (1992), 365–378.
- Heger et al. (2022) Amy K. Heger, Liz B. Marquis, Mihaela Vorvoreanu, Hanna Wallach, and Jennifer Wortman Vaughan. 2022. Understanding Machine Learning Practitioners’ Data Documentation Perceptions, Needs, Challenges, and Desiderata. arXiv (August 24 2022). http://arxiv.org/abs/2206.02923
- Hind et al. (2019) Michael Hind, Stephanie Houde, Jacquelyn Martino, Aleksandra Mojsilovic, David Piorkowski, John Richards, and Kush R. Varshney. 2019. Experiences with Improving the Transparency of AI Models and Services. arXiv (November 11 2019). http://arxiv.org/abs/1911.08293
- Holland et al. (2018) Sarah Holland, Ahmed Hosny, Sarah Newman, Joshua Joseph, and Kasia Chmielinski. 2018. The Dataset Nutrition Label: A Framework To Drive Higher Data Quality Standards.
- Hupont and Gomez (2022) Isabelle Hupont and Emilia Gomez. 2022. Documenting Use Cases in the Affective Computing Domain Using Unified Modeling Language. arXiv (September 19 2022). http://arxiv.org/abs/2209.09666
- Hutchinson et al. (2022) Ben Hutchinson, Negar Rostamzadeh, Christina Greer, Katherine Heller, and Vinodkumar Prabhakaran. 2022. Evaluation Gaps in Machine Learning Practice. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 1859–1876. https://doi.org/10.1145/3531146.3533233
- Jones (2023) Elliot Jones. 2023. Foundation Models: An Explainer. https://www.adalovelaceinstitute.org/resource/foundation-models-explainer/ Accessed: 2024-08-27.
- Liang et al. (2024) Weixin Liang, Nazneen Rajani, Xinyu Yang, Ezinwanne Ozoani, Eric Wu, Yiqun Chen, Daniel Scott Smith, and James Zou. 2024. What’s Documented in AI? Systematic Analysis of 32K AI Model Cards. arXiv (February 7 2024). http://arxiv.org/abs/2402.05160
- Liao et al. (2023) Q. Vera Liao, Hariharan Subramonyam, Jennifer Wang, and Jennifer Wortman Vaughan. 2023. Designerly Understanding: Information Needs for Model Transparency to Support Design Ideation for AI-Powered User Experience. arXiv (February 20 2023). http://arxiv.org/abs/2302.10395
- Liao and Vaughan (2023) Q Vera Liao and Jennifer Wortman Vaughan. 2023. Ai transparency in the age of llms: A human-centered research roadmap. arXiv preprint arXiv:2306.01941 (2023), 5368–5393.
- Madaio et al. (2024) Michael Madaio, Shivani Kapania, Rida Qadri, Ding Wang, Andrew Zaldivar, Remi Denton, and Lauren Wilcox. 2024. Learning about Responsible AI On-The-Job: Learning Pathways, Orientations, and Aspirations. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. ACM, Rio de Janeiro Brazil, 1544–1558. https://doi.org/10.1145/3630106.3658988
- Marone and Van Durme (2023) Marc Marone and Benjamin Van Durme. 2023. Data Portraits: Recording Foundation Model Training Data. arXiv (December 14 2023). http://arxiv.org/abs/2303.03919
- McLennan et al. (2022) Stuart McLennan, Amelia Fiske, Daniel Tigard, Ruth Müller, Sami Haddadin, and Alena Buyx. 2022. Embedded Ethics: A Proposal for Integrating Ethics into the Development of Medical AI. BMC Medical Ethics 23, 1 (December 2022), 6. https://doi.org/10.1186/s12910-022-00746-3
- McMillan-Major et al. (2024) Angelina McMillan-Major, Emily M. Bender, and Batya Friedman. 2024. Data Statements: From Technical Concept to Community Practice. ACM Journal on Responsible Computing 1, 1 (March 31 2024), 1–17. https://doi.org/10.1145/3594737
- McMillan-Major et al. (2021) Angelina McMillan-Major, Salomey Osei, Juan Diego Rodriguez, Pawan Sasanka Ammanamanchi, Sebastian Gehrmann, and Yacine Jernite. 2021. Reusable Templates and Guides For Documenting Datasets and Models for Natural Language Processing and Generation: A Case Study of the HuggingFace and GEM Data and Model Cards. In Proceedings of the 1st Workshop on Natural Language Generation, Evaluation, and Metrics (GEM 2021). 121–135. https://doi.org/10.18653/v1/2021.gem-1.11
- Miceli et al. (2021) Milagros Miceli, Tianling Yang, Laurens Naudts, Martin Schuessler, Diana Serbanescu, and Alex Hanna. 2021. Documenting Computer Vision Datasets: An Invitation to Reflexive Data Practices. In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency. ACM, Virtual Event Canada, 161–172. https://doi.org/10.1145/3442188.3445880
- Micheli et al. (2023) Marina Micheli, Isabelle Hupont, Blagoj Delipetrev, and Josep Soler-Garrido. 2023. The Landscape of Data and AI Documentation Approaches in the European Policy Context. Ethics and Information Technology 25, 4 (October 28 2023), 56. https://doi.org/10.1007/s10676-023-09725-7
- Mitchell et al. (2019) Margaret Mitchell, Simone Wu, Andrew Zaldivar, Parker Barnes, Lucy Vasserman, Ben Hutchinson, Elena Spitzer, et al. 2019. Model Cards for Model Reporting. In Proceedings of the Conference on Fairness, Accountability, and Transparency. 220–229. https://doi.org/10.1145/3287560.3287596
- Mohammad (2022) Saif M. Mohammad. 2022. Ethics Sheets for AI Tasks. arXiv (March 19 2022). http://arxiv.org/abs/2107.01183
- Mökander et al. (2023) Jakob Mökander, Jonas Schuett, Hannah Rose Kirk, and Luciano Floridi. 2023. Auditing Large Language Models: A Three-Layered Approach. AI and Ethics (May 30 2023). https://doi.org/10.1007/s43681-023-00289-2
- Moore et al. (2023) Steven Moore, Q. Vera Liao, and Hariharan Subramonyam. 2023. fAIlureNotes: Supporting Designers in Understanding the Limits of AI Models for Computer Vision Tasks. arXiv (February 22 2023). https://doi.org/10.48550/arXiv.2302.11703
- Narayanan and Kapoor (2024) Arvind Narayanan and Sayash Kapoor. 2024. AI Safety Is Not a Model Property. https://www.aisnakeoil.com/p/ai-safety-is-not-a-model-property
- Nicholas (2024) Gabriel Nicholas. 2024. Grounding AI Policy: Towards Researcher Access to AI Usage Data. https://cdt.org/wp-content/uploads/2024/08/2024-08-12-CDT-Research-Grounding-AI-Policy-report-final.pdf
- Nunes et al. (2022) José Luiz Nunes, Gabriel D. J. Barbosa, Clarisse Sieckenius De Souza, Helio Lopes, and Simone D. J. Barbosa. 2022. Using Model Cards for Ethical Reflection: A Qualitative Exploration. In Proceedings of the 21st Brazilian Symposium on Human Factors in Computing Systems. ACM, Diamantina Brazil, 1–11. https://doi.org/10.1145/3554364.3559117
- OpenAI (2023) OpenAI. 2023. GPT-4V(Ision) System Card. https://openai.com/index/gpt-4v-system-card/
- Palinkas et al. (2015) Lawrence A Palinkas, Sarah M Horwitz, Carla A Green, Jennifer P Wisdom, Naihua Duan, and Kimberly Hoagwood. 2015. Purposeful sampling for qualitative data collection and analysis in mixed method implementation research. Administration and policy in mental health and mental health services research 42 (2015), 533–544.
- Papakyriakopoulos et al. (2023) Orestis Papakyriakopoulos, Anna Seo Gyeong Choi, William Thong, Dora Zhao, Jerone Andrews, Rebecca Bourke, Alice Xiang, and Allison Koenecke. 2023. Augmented Datasheets for Speech Datasets and Ethical Decision-Making. In 2023 ACM Conference on Fairness, Accountability, and Transparency. ACM, Chicago IL USA, 881–904. https://doi.org/10.1145/3593013.3594049
- Pepe et al. (2024) Federica Pepe, Vittoria Nardone, Antonio Mastropaolo, Gerardo Canfora, and Gabriele Bavota. 2024. How Do Hugging Face Models Document Datasets, Bias, and Licenses? An Empirical Study. (2024).
- Phillips-Wren and Adya (2020) Gloria Phillips-Wren and Monica Adya. 2020. Decision Making under Stress: The Role of Information Overload, Time Pressure, Complexity, and Uncertainty. Journal of Decision Systems 29, sup1 (August 18 2020), 213–225. https://doi.org/10.1080/12460125.2020.1768680
- Procope et al. (2022) Chavez Procope, Adeel Cheema, David Adkins, Bilal Alsallakh, Emily McReynolds, Grace Pehl, Erin Wang, and Polina Zvyagina. 2022. System-Level Transparency of Machine Learning. (February 22 2022). https://ai.meta.com/research/publications/system-level-transparency-of-machine-learning/
- Pushkarna et al. (2022) Mahima Pushkarna, Andrew Zaldivar, and Oddur Kjartansson. 2022. Data Cards: Purposeful and Transparent Dataset Documentation for Responsible AI. arXiv (April 3 2022). http://arxiv.org/abs/2204.01075
- Raji and Yang (2020) Inioluwa Deborah Raji and Jingying Yang. 2020. ABOUT ML: Annotation and Benchmarking on Understanding and Transparency of Machine Learning Lifecycles. arXiv (January 7 2020). http://arxiv.org/abs/1912.06166
- Reid and Williams (2023) Kathy Reid and Elizabeth T. Williams. 2023. Right the Docs: Characterising Voice Dataset Documentation Practices Used in Machine Learning. arXiv (March 19 2023). http://arxiv.org/abs/2303.10721
- Richards et al. (2020) John Richards, David Piorkowski, Michael Hind, Stephanie Houde, and Aleksandra Mojsilović. 2020. A Methodology for Creating AI FactSheets. arXiv (June 27 2020). https://doi.org/10.48550/arXiv.2006.13796
- Roman et al. (2023) Anthony Cintron Roman, Jennifer Wortman Vaughan, Valerie See, Steph Ballard, Nicolas Schifano, Jehu Torres, Caleb Robinson, and Juan M. Lavista Ferres. 2023. Open Datasheets: Machine-Readable Documentation for Open Datasets and Responsible AI Assessments. arXiv (December 11 2023). http://arxiv.org/abs/2312.06153
- Rostamzadeh et al. (2022) Negar Rostamzadeh, Diana Mincu, Subhrajit Roy, Andrew Smart, Lauren Wilcox, Mahima Pushkarna, Jessica Schrouff, Razvan Amironesei, Nyalleng Moorosi, and Katherine Heller. 2022. Healthsheet: Development of a Transparency Artifact for Health Datasets. arXiv (February 25 2022). http://arxiv.org/abs/2202.13028
- Shen et al. (2022) Hong Shen, Leijie Wang, Wesley H. Deng, Ciell Brusse, Ronald Velgersdijk, and Haiyi Zhu. 2022. The Model Card Authoring Toolkit: Toward Community-Centered, Deliberation-Driven AI Design. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 440–451. https://doi.org/10.1145/3531146.3533110
- Shilton (2013) Katie Shilton. 2013. Values Levers: Building Ethics into Design. Science, Technology, & Human Values 38, 3 (May 2013), 374–397. https://doi.org/10.1177/0162243912436985
- Shimorina and Belz (2021) Anastasia Shimorina and Anya Belz. 2021. The Human Evaluation Datasheet 1.0: A Template for Recording Details of Human Evaluation Experiments in NLP. arXiv (March 17 2021). http://arxiv.org/abs/2103.09710
- Smart et al. (2024) Andrew Smart, Shazeda Ahmed, Jacob Metcalf, Atoosa Kasirzadeh, Luca Belli, Shalaleh Rismani, Roel Dobbe, et al. 2024. What Is Sociotechnical AI Safety? What Do We Want It To Be? A FAccT Community Workshop.
- Soh (2021) Jerrold Soh. 2021. Building Legal Datasets. arXiv (November 3 2021). http://arxiv.org/abs/2111.02034
- Sokol and Flach (2020) Kacper Sokol and Peter Flach. 2020. Explainability Fact Sheets: A Framework for Systematic Assessment of Explainable Approaches. In Proceedings of the 2020 Conference on Fairness, Accountability, and Transparency. ACM, Barcelona Spain, 56–67. https://doi.org/10.1145/3351095.3372870
- Srinivasan et al. (2021) Ramya Srinivasan, Emily Denton, Jordan Famularo, Negar Rostamzadeh, Fernando Diaz, and Beth Coleman. 2021. Artsheets for Art Datasets.
- Stoyanovich and Howe (2019) Julia Stoyanovich and Bill Howe. 2019. Nutritional Labels for Data and Models. Bulletin of the IEEE Computer Society Technical Committee on Data Engineering (2019).
- Strauss and Corbin (1990) Anselm Strauss and Juliet Corbin. 1990. Basics of qualitative research. Sage Publications.
- Subramaniam et al. (2023) Pranav Subramaniam, Yintong Ma, Chi Li, Ipsita Mohanty, and Raul Castro Fernandez. 2023. Comprehensive and Comprehensible Data Catalogs: The What, Who, Where, When, Why, and How of Metadata Management. arXiv (February 1 2023). http://arxiv.org/abs/2103.07532
- Sun et al. (2019) Chenkai Sun, Abolfazl Asudeh, H. V. Jagadish, Bill Howe, and Julia Stoyanovich. 2019. MithraLabel: Flexible Dataset Nutritional Labels for Responsible Data Science. In Proceedings of the 28th ACM International Conference on Information and Knowledge Management. ACM, Beijing China, 2893–2896. https://doi.org/10.1145/3357384.3357853
- Tagliabue et al. (2021) Jacopo Tagliabue, Ville Tuulos, Ciro Greco, and Valay Dave. 2021. DAG Card Is the New Model Card. arXiv (November 20 2021). http://arxiv.org/abs/2110.13601
- Tavory and Timmermans (2014) Iddo Tavory and Stefan Timmermans. 2014. Abductive analysis: Theorizing Qualitative Research. University of Chicago Press.
- Timmermans and Tavory (2012) Stefan Timmermans and Iddo Tavory. 2012. Theory construction in qualitative research: From grounded theory to abductive analysis. Sociological Theory 30, 3 (2012), 167–186.
- Touvron et al. (2023) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, et al. 2023. Llama 2: Open Foundation and Fine-Tuned Chat Models. https://ai.meta.com/research/publications/llama-2-open-foundation-and-fine-tuned-chat-models/
- Vallor et al. (2018) Shannon Vallor, Irina Raicu, and Brian Green. 2018. Overview of Ethics in Tech Practice. https://www.scu.edu/ethics-in-technology-practice/overview-of-ethics-in-tech-practice/
- Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N. Gomez, Lukasz Kaiser, and Illia Polosukhin. 2017. Attention Is All You Need. https://doi.org/10.48550/arXiv.1706.03762
- Winecoff and Watkins (2022) Amy A. Winecoff and Elizabeth Anne Watkins. 2022. Artificial Concepts of Artificial Intelligence: Institutional Compliance and Resistance in AI Startups. In Proceedings of the 2022 AAAI/ACM Conference on AI, Ethics, and Society. ACM, Oxford United Kingdom, 788–799. https://doi.org/10.1145/3514094.3534138
- Yang et al. (2018) Ke Yang, Julia Stoyanovich, Abolfazl Asudeh, Bill Howe, H. V. Jagadish, and Gerome Miklau. 2018. A Nutritional Label for Rankings. In Proceedings of the 2018 International Conference on Management of Data. ACM, Houston TX USA, 1773–1776. https://doi.org/10.1145/3183713.3193568
- Zheng et al. (2022) Xinyi Zheng, Ryan A. Rossi, Nesreen Ahmed, and Dominik Moritz. 2022. Network Report: A Structured Description for Network Datasets. arXiv (June 7 2022). http://arxiv.org/abs/2206.03635

### v2

@@
 Although there are many proposals for documenting AI data, models, systems, and methods, the ways these practices enhance governance as well as the challenges practitioners and organizations face with documentation remain underexplored.
-In this paper, we analyze 37 proposed documentation frameworks and 21 empirical studies evaluating their use.
-We identify potential hypotheses about how documentation can strengthen governance—such as informing stakeholders about AI risks and usage, fostering collaboration, encouraging ethical reflection, and reinforcing best practices.
-However, empirical evidence shows that practitioners often encounter obstacles that prevent documentation from achieving these goals.
+In this paper, we analyze 37 proposed documentation frameworks and 22 empirical studies evaluating their use.
+We identify several pathways or ”theories of change” through which documentation can enhance governance, including informing stakeholders about AI risks and applications, facilitating collaboration, encouraging ethical deliberation, and supporting best practices.
+However, empirical findings reveal significant challenges for practitioners, such as insufficient incentives and resources, structural and organizational communication barriers, interpersonal and organizational constraints to ethical action, and poor integration with existing workflows.
+These challenges often hinder the realization of the possible benefits of documentation.
 We also highlight key considerations for organizations when designing documentation, such as determining the appropriate level of detail and balancing automation in the process.
-Finally, we offer recommendations for further research and for implementing effective documentation practices in real-world contexts.
+We conclude by discussing how future research can expand on our findings such as by exploring documentation approaches that support governance of general-purpose models and how multiple transparency and documentation methods can collectively improve governance outcomes.
 1. Introduction
-Transparency in AI is widely regarded as essential for calibrating trust and supporting accountability (Liao and Vaughan, 2023).
+Transparency in AI is widely regarded as essential for calibrating trust and supporting accountability (Liao and Vaughan, 2024).
 By providing stakeholders with a clear view of an AI system’s composition, operation, and development process, transparency allows for informed oversight and critical evaluation.
 When AI systems are thoroughly documented, these documentation artifacts offer invaluable insights into important system features such as training data, algorithms, and risk management strategies.
-This information can support two broad groups of external stakeholders.
+This information can support several groups of external stakeholders.
 Documentation can help downstream deployers and users understand how the system functions and what risks it conveys.
@@
 Yet the importance of documentation extends beyond external accountability.111Though the concepts of documentation and transparency are often blurred, complete documentation about an AI system does not guarantee that information will be made available to interested stakeholders, nor do transparency artifacts always reflect the full or most relevant details of an AI system even though such details may have been captured in internal documents.
-We distinguish between the concepts in this report in order to provide greater clarity about the presumptive goals of documentation and an exploration of what practices may more likely support those objectives Internally, documentation serves as a critical tool for managing AI systems throughout their lifecycle for a wide range of stakeholders.
+We distinguish between the concepts in this report in order to provide greater clarity about the presumptive goals of documentation and an exploration of what practices may more likely support those objectives.
+Internally, documentation serves as a critical tool for managing AI systems throughout their lifecycle for a wide range of stakeholders.
 For example, documentation could help downstream technical practitioners understand the strengths, limitations, and risks of training a model on a given dataset.
@@
 Or the process of producing documentation might lead practitioners to make different decisions about how to sufficiently mitigate potential harms.
-As such, documentation is a foundational building block of effective AI governance.
+Robust documentation can improve the efficiency and quality of other important AI governance practices and processes, such as internal and third-party auditing, harm identification, risk triaging, and compliance functions, and can help practitioners better implement risk management practices, such as system evaluation and risk mitigation approaches.
+In this way, documentation serves as a foundational building block of a multifaceted AI governance approach, both providing important inputs to other governance processes and creating a scaffold on which additional governance practices can build over time.
 Despite its clear benefits, creating effective documentation is a complex task.
@@
 Effective documentation design requires balancing the needs of both the documentation process and the resulting artifacts, ensuring that they collectively support the organization’s governance objectives.
-In this paper, we explore how documentation processes and artifacts can best support AI governance and risk management goals.
-Our research draws from a qualitative analysis of academic and gray literature,222Gray literature refers to publications that have not been peer-reviewed, but that present detailed theories or research findings including 37 proposed documentation frameworks and 21 papers with relevant empirical findings.
-From this analysis, we identify four mechanisms through which documentation is hypothesized to improve AI governance:
-Informing stakeholders about responsible use,
-Facilitating collaboration on AI risks,
-Encouraging ethical deliberation, and
-Improving overall governance and development practices.
-We also examine the practical barriers that may hinder the effectiveness of these mechanisms and the considerations or trade-offs organizations face when designing and implementing documentation practices.
-These considerations include balancing customization with standardization, deciding between single or multiple forms of documentation, determining the extent of detail to include, choosing the appropriate level of automation to employ in the documentation process, and selecting between static or interactive formats.
-Based on our analysis, we offer recommendations to help organizations develop documentation strategies that align with their governance goals and stakeholder needs.
+In this paper, we explore how documentation processes and artifacts can serve as one critical building block in supporting AI governance and risk management goals.
+Our research draws from a qualitative analysis of academic and gray literature,222Gray literature refers to publications that have not been peer-reviewed, but that present detailed theories or research findings including 37 proposed documentation frameworks and 22 papers with relevant empirical findings.
+From this analysis, we identify four ”theories of change” or pathways through which documentation could improve AI governance, including informing stakeholders about responsible use, facilitating collaboration on AI risks, encouraging ethical deliberation, and improving overall governance and development practices.
+Our analysis also identified several challenges that might prevent documentation from effectively supporting governance aims.
+These include insufficient incentivization and resourcing for documentation, structural and organizational barriers to effective collaboration, interpersonal and organizational constraints to ethical action, and poor integration with practitioners’ day-to-day workflows and processes.
+We also explored considerations or trade-offs organizations face when designing and implementing documentation practices to balance competing priorities or documentation goals.
+These considerations include balancing customization with standardization, deciding between single or multiple forms of documentation to address different audiences, determining the extent of detail to include, choosing the appropriate level of automation to employ in the documentation process, and selecting between static or interactive formats.
 The primary contribution of our research is that our synthesis identifies patterns, gaps, and emerging trends across individual publications about documentation approaches and evaluations.
@@
 As such, our research can serves as a valuable resource for both researchers and practitioners seeking to refine and enhance their approaches to AI documentation in applied AI development contexts.
+We conclude with a discussion of how future research can build on or expand upon the themes we explore throughout the paper.
 2. What organizations could document about AI systems
@@
 Documenting aspects such as the model’s architecture, training procedures, intended use cases, and evaluation methods enables practitioners to assess potential risks and responsibly adapt models to specific contexts.
-For models developed in stages—such as those pretrained on large datasets and later fine-tuned—documenting each stage is essential.
+For models developed in stages—such as those pretrained on large datasets and later fine-tuned—documenting each stage is essential (Gilbert et al., 2023).
 This helps downstream users understand the development pipeline and its implications for their intended use.
@@
 The interaction among these components significantly affects the system’s overall performance and risk profile — especially because the properties or behavior of individual components may change when integrated into a system (Smart et al., 2024).
-Recognizing that AI systems are more than just the sum of their parts, researchers have suggested that organizations should document not only each individual component but also the system as a whole (Arnold et al., 2019; Baracaldo et al., 2022; Brajovic et al., 2023; Green et al., 2021; Procope et al., 2022; Yang et al., 2018; Blasch et al., 2021).
-System documentation could provide an overview, including objectives, inputs, outputs, and a diagram illustrating how the components interact.
+Recognizing that AI systems are more than just the sum of their parts, researchers have suggested that organizations should document not only each individual component but also the system as a whole (Arnold et al., 2019; Baracaldo et al., 2022; Brajovic et al., 2023; Procope et al., 2022; Yang et al., 2018; Blasch et al., 2021).
+System documentation could provide an overview, including objectives, inputs, outputs, and a diagram illustrating how the components interact (Procope et al., 2022).
 This comprehensive approach ensures that practitioners have a clear understanding of the system’s behavior in its entirety.
@@
 Documentation producers are those who actively contribute to the generation of documentation artifacts, while documentation consumers are those who read and use it.
-While this grouping simplifies some important distinctions, it serves as a helpful heuristic for understanding the opportunities and tensions between groups with both shared and distinct goals.
+This grouping simplifies some important distinctions.333Practitioners sometimes create documentation for their own use, while in other instances, they produce one type of documentation for others and rely on a different type created by someone else.
+For example, a compliance team might use documentation from a technical team to understand how a system was developed and functions in order to assess compliance concerns.
+At the same time, the compliance team might generate documentation for technical teams, outlining the limitations that apply to the use of a system or component.
+In this way, the same stakeholder could be both a consumer and a producer of documentation, even for the same system or component.
+Still, distinguishing between consumers and producers serves as a helpful heuristic for understanding the opportunities and tensions between groups with both shared and distinct goals.
 We consider both producers and consumers and their interplay in our analysis.
 We also note that ”AI documentation” could refer to either the process of producing documentation or the resulting artifacts.
-While documentation artifacts inform stakeholders about the responsible use of AI, the process of creating these documents can also institutionalize best practices and foster a culture of risk management..
+While documentation artifacts can inform stakeholders about the responsible use of AI, the process of creating these documents can also institutionalize best practices and foster a culture of risk management.
 Therefore, when considering how documentation can improve governance outcomes, we address both artifacts and processes.
@@
 3.1. Publication Sampling
-We employed a purposeful sampling approach (Palinkas et al., 2015) to identify proposed documentation frameworks and relevant empirical studies.
-Given our specific interest in how AI documentation can support governance, our use of a purposeful sampling approach allowed us to focus on proposed frameworks and empirical studies that are directly relevant to our research questions.
-It also afforded us the flexibility to iteratively refine our sample of papers based on emerging findings.
+We employed a purposeful sampling approach (Palinkas et al., 2015; Suri, 2011) to identify publications.
+This method was particularly well-suited to our study because it allowed us to target information-rich frameworks and empirical studies clearly aligned with our research goals of understanding the functions documentation plays in supporting governance and the challenges associated with using documentation as a foundational governance tool.
+By prioritizing publications that offered insights into documentation frameworks, evaluation methods, and practical applications, we ensured extensive coverage of existing theories and practices while maintaining direct relevance to our research questions.
+Furthermore, our iterative sampling process enabled us to refine our dataset as our analysis progressed, incorporating studies that contributed meaningful data or theoretical perspectives.
+Our approach contrasts with systematic sampling, which establishes strict inclusion and exclusion criteria beforehand.
+While systematic sampling might have offered more exhaustive coverage, it likely would have included less relevant papers that did not align as closely with our objectives and would not have allowed us to expand our sample in response to emerging findings.
+We opted to include both proposed frameworks and empirical evaluations in our analysis rather than limiting our focus to empirical studies.
+This decision stemmed from the recognition that the goals of a documentation framework and its real-world implementation can often diverge.
+For instance, a framework that generates high-quality documentation in a research context may fall short when applied in practical settings.
+By examining both the intended goals of proposed frameworks and the empirical evidence, we aimed to better understand and highlight the gap between these two dimensions, informing future designs and evaluations.
 We identified an initial seed sample of publications from a systematic review focused on AI documentation methods relevant to EU regulation (Micheli et al., 2023).
-We cross-referenced this initial sample against the references used by the Partnership on AI in developing their ABOUT ML framework,333https://partnershiponai.org/workstream/about-ml/ which seeks to provide technology organizations with guidance on what to document about their AI systems.
+We cross-referenced this initial sample against the references used by the Partnership on AI in developing their ABOUT ML framework,444https://partnershiponai.org/workstream/about-ml/ which seeks to provide technology organizations with guidance on what to document about their AI systems.
 We chose these initial sources because they allowed us to focus on publications that aimed to establish best practices for industry documentation and address policy initiatives geared toward institutionalizing these practices.
-We reviewed the cited references within each of these initial works for additional proposed documentation frameworks and empirical studies related to documentation.
-We also consulted with academic researchers and industry practitioners with experience producing and using documentation to identify additional frameworks and studies.
-We excluded papers that centered on tools for implementing documentation (e.g., automated documentation code libraries) or that were not specifically focused on AI system documentation, such as fairness checklists.
+We reviewed the references of these publications and consulted with subject matter experts in industry and academia with experience producing or using documentation to identify additional publications.
 This process yielded a sample of 37 proposed frameworks, eight of which included empirical evidence related to their implementation.
-Because of the relative dearth of empirical findings within publications proposing approaches, we further expanded our sample by searching for references within these 37 approaches that empirically evaluated documentation methods.
-We also conducted searches in the archives of the ACM Conference on Human Factors in Computing Systems and the ACM Conference on Computer Supported Cooperative Work using terms such as ”datasheet,” ”model card,” ”AI documentation,” ”model documentation,” ”data card,” and ”data documentation.” Together, these methods allowed us to identify 13 additional publications with empirical findings.
+To expand our sample of empirical findings, we again consulted with subject matter experts and reviewed the citations of the frameworks for references to further empirical work.
+We also conducted searches in the archives of the ACM Conference on Human Factors in Computing Systems and the ACM Conference on Computer Supported Cooperative Work using terms such as ”datasheet,” ”model card,” ”AI documentation,” ”model documentation,” ”data card,” and ”data documentation.” Together, these methods allowed us to identify 14 additional publications with empirical findings.
+Our sampling concluded upon reaching theoretical saturation, where further sampling no longer yielded new themes, supporting the adequacy and richness of our chosen publications.
 Recognizing the importance of applied perspectives, we included non-peer-reviewed papers, such as technical whitepapers published by industry actors, acknowledging that they often provide insights not found in academic literature.
-While our sampling approach was not exhaustive, we reached theoretical saturation after analyzing the selected publications, as no new themes emerged.
-Therefore, we concluded our sampling at this point.
+To maintain a targeted focus on publications examining the governance role of documentation rather than the technical aspects of specific tools, we excluded papers discussing enabling technologies (e.g., automated documentation libraries or supporting software) or those addressing broader topics like fairness checklists not specific to AI documentation.
 3.2. Data Analysis
-Our data analysis followed an abductive approach (Timmermans and Tavory, 2012; Tavory and Timmermans, 2014), iterating between inductively derived codes, as in grounded theory methods (Strauss and Corbin, 1990; Charmaz, 2014), and codes based on the theoretical motivations of our research.
-The first author conducted the initial coding by applying descriptive codes aligned with our research goals, such as identifying how AI documentation enhances internal governance and the challenges of using documentation as a governance tool.
-After discussing the results with the second author, we refined these codes into higher-order categories during the axial coding phase.
-For instance, codes like ”identify intended purpose” and ”specify out-of-scope uses” were grouped under the category ”inform downstream documentation consumers.”
-In the final thematic coding phase, we grouped axial codes into broader theoretical mechanisms through which AI documentation could support robust governance, the challenges in achieving these impacts, and the design trade-offs faced by organizations.
+We employed an abductive approach for our data analysis (Tavory and Timmermans, 2014; Timmermans and Tavory, 2012), a qualitative method that involves alternating between theoretical concepts identified deductively and codes generated inductively through data review, similar to grounded theory methods.
+(Charmaz, 2014; Strauss and Corbin, 1990; Glaser, 1978; Glaser et al., 1968).
+Guided by this approach, we initially applied codes aligned with our objectives of understanding how documentation can support robust AI governance and identifying barriers to its effectiveness.
+At the same time, we remained open to adjusting our coding focus based on insights gained through inductive analysis of our publication sample.
+Our analysis was refined iteratively through ongoing discussions and repeated reviews of the sample publications in alignment with these objectives.
+In the initial coding phase, the first author applied descriptive codes to papers that aligned with our research objectives.
+These codes included terms like “discoverability,” “standardization,” “intended use,” “incentivization,” “documentation errors.” During the review of these descriptive codes, both authors observed emerging themes related to both theories of change and challenges.
+For example, we noted that both the codes ”identify intended purpose” and ”specify out of scope uses” both pertained to informing downstream documentation consumers.
+Following this discussion, we organized our descriptive codes into four theory of change categories: informing downstream stakeholders, facilitating collaboration and communication, promoting ethical deliberation, and supporting governance and development best practices.
+We grouped codes related to challenges into four overarching categories: insufficient incentives and resources, structural and organizational obstacles to effective communication, limited ethical awareness and barriers to ethical action, and inadequate integration with existing workflows and organizational practices.
+We note that in some cases, challenges are more relevant to specific theories of change, and in other cases, challenges are more generally applicable.
+We discuss these points of connection in the results.
+The inductive aspects of our analysis highlighted that authors of documentation publications made varied design choices aligned with the goals of their proposed frameworks.
+Consequently, we expanded our focus to examine the design tensions and considerations that surfaced within and across these papers.
+To capture these dynamics, we grouped descriptive codes related to design choices, particularly those that appeared to exist in tension with one another.
+These considerations included the balance between customization and standardization in addressing specific components or use cases, the extent of audience-specific tailoring in documentation artifacts, the level of interactivity provided, the amount of detail included, and the degree of automation in the documentation process.
+3.3. Positionality
+The authors of this study work at a civil society organization specializing in technology policy.
+Previously, the first author developed machine learning models, and the second author focused on AI governance and policy within technology companies.
+Both authors have experience conducting research in human-centered and responsible AI and have advised government and industry on implementing risk management practices for AI systems.
+Our academic backgrounds in psychology and public policy inform our understanding of AI as a sociotechnical system and AI governance as a sociotechnical process.
+These perspectives inevitably shaped our selection of publications, the themes we chose to explore, and our synthesis of the results.
+Given that our analysis approach is interpretive, researchers with different perspectives and backgrounds may have arrived at different conclusions when analyzing the same dataset or found different publications more meaningful to the overall research aims.
 4. Results
@@
 When a framework did not fit neatly into one of these categories, we assigned it to the most appropriate category or categories based on its primary focus.
-We also categorized the type of evaluation each framework or independent empirical research study employed.
+We also categorized the type of evaluation each framework or empirical research study employed.
 Frameworks employing a ”feasibility analysis” are those where the framework’s authors or another group applied the framework to create documentation for a hypothetical or actual dataset, model, system, or method.
@@
 In some instances, framework authors mentioned consulting relevant stakeholders during the design or refinement of their framework.
-However, if these consultations are only briefly mentioned or not elaborated upon, we do not classify the work as including a practitioner study.
+However, if these consultations were only briefly mentioned, we do not classify the work as including a practitioner study.
 | Num. | Author | Year | Framework Type | Evaluation Type |
 | 1 | Adkins et al. | 2022 | Method, process, or task | Feasibility analysis |
-| 2 | Arnold et al. | 2019 | System | Feasibility analysis |
-| 3 | Baracaldo et al. | 2022 | System | Feasibility analysis |
-| 4 | Bender & Friedman | 2018 | Data | Feasibility analysis |
-| 5 | Bhat et al. | 2023 | Evaluation only | Practitioner lab study |
-| 6 | Blasch et al. | 2020 | System | Feasibility analysis |
-| 7 | Boyd | 2021 | Evaluation only | Practitioner lab study |
-| 8 | Brajovic et al. | 2023 | System | Feasibility analysis |
-| 9 | Chang & Custis | 2022 | Evaluation only | Practitioner real-world study |
-| 10 | Chmielinski et al. | 2020 | Data | Feasibility analysis |
-| 11 | Chmielinski et al. | 2024 | Method, process, or task | None |
-| 12 | Crisan et al. | 2022 | Model | Feasibility analysis, Practitioner lab study |
-| 13 | Díaz et al. | 2023 | Data | Feasibility analysis |
-| 14 | Gebru et al. | 2021 | Data | Feasibility analysis |
-| 15 | Geiger et al. | 2020 | Evaluation only | Artifact study |
-| 16 | Gilbert et al. | 2023 | System | Feasibility analysis |
-| 17 | Heger et al. | 2022 | Evaluation only | Practitioner real-world study |
-| 18 | Hind et al. | 2019 | Evaluation only | Practitioner real-world study |
-| 19 | Holland et al. | 2018 | Data | Feasibility analysis |
-| 20 | Hupont & Gomez | 2022 | Method, process, or task | Feasibility analysis |
-| 21 | Liang et al. | 2024 | Evaluation only | Artifact study |
-| 22 | Liao et al. | 2023 | Evaluation only | Practitioner lab study |
-| 23 | Marone & Van Durme | 2023 | Data | Feasibility analysis |
-| 24 | McMillan-Major et al. | 2021 | Data, Model | Feasibility analysis |
-| 25 | McMillan-Major, Bender, & Friedman | 2024 | Data | Practitioner lab study |
-| 26 | Miceli et al. | 2021 | Evaluation only | Practitioner real-world study |
-| 27 | Mitchell et al. | 2019 | Model | Feasibility analysis |
-| 28 | Mohammad | 2022 | Method, process, or task | Feasibility analysis |
-| 29 | Moore, LIao, & Subramonyam | 2023 | Evaluation only | Practitioner lab study |
-| 30 | Nunes et al. | 2022 | Evaluation only | Practitioner lab study |
-| 31 | Papakyriakopoulos et al. | 2023 | Data | Feasibility analysis |
-| 32 | Pepe et al. | 2024 | Evaluation only | Artifact study |
-| 33 | Procope et al. | 2022 | System | Feasibility analysis |
-| 34 | Pushkarna et al. | 2022 | Data | Practitioner real-world study |
-| 35 | Raji & Yang | 2020 | System | None |
-| 36 | Reid & Williams | 2023 | Evaluation only | Practitioner real-world study, Artifact study |
-| 37 | Richards et al. | 2020 | System | Feasibility analysis |
-| 38 | Roman et al. | 2023 | Data | Feasibility analysis |
-| 39 | Rostamzadeh et al. | 2022 | Data | Practitioner lab study |
-| 40 | Shen et al. | 2022 | Model | Practitioner lab study |
-| 41 | Shimorina & Belz | 2021 | Method, process, or task | None |
-| 42 | Soh | 2021 | Data | None |
-| 43 | Sokol & Flach | 2020 | Method, process, or task | Feasibility analysis |
-| 44 | Srinivasan et al. | 2021 | Data | Feasibility analysis |
-| 45 | Stoyanovich & Howe | 2019 | Data, Model | Feasibility analysis |
-| 46 | Subramaniam et al. | 2023 | Data | Practitioner lab study |
-| 47 | Sun et al. | 2019 | Data | Feasibility analysis |
-| 48 | Tagliabue et al. | 2021 | Method, process, or task | Feasibility analysis |
-| 49 | Yang et al. | 2018 | Method, process, or task | Feasibility analysis |
-| 50 | Zheng et al., | 2022 | Data | Practitioner lab study |
-4.2. Potential Impacts & Challenges
-While empirical evidence directly linking documentation artifacts or processes to improved governance outcomes is limited, the theories proposed in existing documentation frameworks offer compelling hypotheses about how documentation can positively impact governance.
-Empirical studies involving practitioners, alongside analyses of publicly available documentation artifacts, provide an initial glimpse into the factors that contribute to accurate, high-quality, and comprehensive documentation, as well as the reasons why documentation efforts may sometimes fall short of their goals.
-Our analysis identified four possible ways in which documentation could improve governance: informing downstream users about system development and associated risks, encouraging ethical reflection among practitioners, facilitating communication among stakeholders, and enhancing AI development and governance overall.
-In the following sections, we explore each of these hypothesized impacts in detail and examine the challenges real-world organizations face in realizing these benefits.
-4.2.1. Potential impact #1: Documentation could inform downstream stakeholders about effective and responsible use.
-The most straightforward potential impact of documentation is that it can provide essential context, such as the original motivation for developing a dataset, model, or system, their intended use cases, and information on system properties necessary for effective and responsible deployment.
-Hypotheses.
-Documentation could inform downstream stakeholders by clearly articulating the intended uses of data, models, and systems and provide guidance on responsible deployment.
-Practitioners typically develop systems with specific use cases in mind, and other development teams need to understand these intended use cases to assess alignment with their goals.
+| 2 | Ahlawat, Winecoff, & Mayer | 2024 | Evaluation only | Practitioner real-world study |
+| 3 | Arnold et al. | 2019 | System | Feasibility analysis |
+| 4 | Baracaldo et al. | 2022 | System | Feasibility analysis |
+| 5 | Bender & Friedman | 2018 | Data | Feasibility analysis |
+| 6 | Bhat et al. | 2023 | Evaluation only | Practitioner lab study |
+| 7 | Blasch et al. | 2020 | System | Feasibility analysis |
+| 8 | Boyd | 2021 | Evaluation only | Practitioner lab study |
+| 9 | Brajovic et al. | 2023 | System | Feasibility analysis |
+| 10 | Chang & Custis | 2022 | Evaluation only | Practitioner real-world study |
+| 11 | Chmielinski et al. | 2020 | Data | Feasibility analysis |
+| 12 | Chmielinski et al. | 2024 | Method, process, or task | None |
+| 13 | Crisan et al. | 2022 | Model | Practitioner lab study |
+| 14 | Díaz et al. | 2023 | Data | Feasibility analysis |
+| 15 | Gebru et al. | 2021 | Data | Feasibility analysis |
+| 16 | Geiger et al. | 2020 | Evaluation only | Artifact study |
+| 17 | Gilbert et al. | 2023 | System | Feasibility analysis |
+| 18 | Heger et al. | 2022 | Evaluation only | Practitioner real-world study |
+| 19 | Hind et al. | 2019 | Evaluation only | Practitioner real-world study |
+| 20 | Holland et al. | 2018 | Data | Feasibility analysis |
+| 21 | Hupont & Gomez | 2022 | Method, process, or task | Feasibility analysis |
+| 22 | Liang et al. | 2024 | Evaluation only | Artifact study |
+| 23 | Liao et al. | 2023 | Evaluation only | Practitioner lab study |
+| 24 | Marone & Van Durme | 2023 | Data | Feasibility analysis |
+| 25 | McMillan-Major et al. | 2021 | Data, Model | Feasibility analysis |
+| 26 | McMillan-Major, Bender, & Friedman | 2024 | Data | Practitioner lab study |
+| 27 | Miceli et al. | 2021 | Evaluation only | Practitioner real-world study |
+| 28 | Mitchell et al. | 2019 | Model | Feasibility analysis |
+| 29 | Mohammad | 2022 | Method, process, or task | Feasibility analysis |
+| 30 | Moore, LIao, & Subramonyam | 2023 | Evaluation only | Practitioner lab study |
+| 31 | Nunes et al. | 2022 | Evaluation only | Practitioner lab study |
+| 32 | Papakyriakopoulos et al. | 2023 | Data | Feasibility analysis |
+| 33 | Pepe et al. | 2024 | Evaluation only | Artifact study |
+| 34 | Procope et al. | 2022 | System | Feasibility analysis |
+| 35 | Pushkarna et al. | 2022 | Data | Practitioner real-world study |
+| 36 | Raji & Yang | 2020 | System | None |
+| 37 | Reid & Williams | 2023 | Evaluation only | Practitioner real-world study, Artifact study |
+| 38 | Richards et al. | 2020 | System | Feasibility analysis |
+| 39 | Roman et al. | 2023 | Data | Feasibility analysis |
+| 40 | Rostamzadeh et al. | 2022 | Data | Practitioner lab study |
+| 41 | Shen et al. | 2022 | Model | Practitioner lab study |
+| 42 | Shimorina & Belz | 2021 | Method, process, or task | None |
+| 43 | Soh | 2021 | Data | None |
+| 44 | Sokol & Flach | 2020 | Method, process, or task | Feasibility analysis |
+| 45 | Srinivasan et al. | 2021 | Data | Feasibility analysis |
+| 46 | Stoyanovich & Howe | 2019 | Data, Model | Feasibility analysis |
+| 47 | Subramaniam et al. | 2023 | Data | Practitioner lab study |
+| 48 | Sun et al. | 2019 | Data | Feasibility analysis |
+| 49 | Tagliabue et al. | 2021 | Method, process, or task | Feasibility analysis |
+| 50 | Yang et al. | 2018 | Method, process, or task | Feasibility analysis |
+| 51 | Zheng et al., | 2022 | Data | Practitioner lab study |
+4.2. Theories of change
+While empirical evidence directly linking documentation artifacts or processes to improved governance outcomes is limited, proposed documentation frameworks offer compelling theories of change about how documentation can positively a multifaceted impact governance approach.
+Our analysis revealed four such theories of change: informing downstream users about system development and associated risks, encouraging ethical reflection among practitioners, facilitating communication among stakeholders, and enhancing AI development and governance overall.
+4.2.1. Theory of change #1: Documentation could inform downstream stakeholders about effective and responsible use.
+Documentation can provide essential background context on the system or component’s development, informing appropriate integration of data, models, methods, or systems into technologies.
+For example, practitioners typically develop systems with specific use cases in mind, and other development teams need to understand these intended use cases to assess alignment with their goals.
 If aligned, teams can then modify the systems to meet their current objectives.
 Practitioners argue documentation can also clarify underlying assumptions, such as data collection methods and representativeness, which are critical for assessing the suitability of a dataset for specific purposes (Reid and Williams, 2023; Rostamzadeh et al., 2022; Zheng et al., 2022).
-Also, it can help identify potential deployment issues, such as mismatches between the original training data and the intended deployment context (Boyd, 2021).
-Moreover, documentation could assist downstream practitioners in making implementation decisions that support effective and responsible product development.
+Also, documentation on background context can help identify potential issues prior to deployment, such as mismatches between the original training data and downstream practitioners’ intended deployment setting (Boyd, 2021).
+Moreover, documentation could provide information that helps downstream practitioners make sound implementation decisions.
 For example, documentation could help practitioners find and compare system components aligned with their goals, including alternatives to AI approaches (Heger et al., 2022; Reid and Williams, 2023; Shen et al., 2022; Zheng et al., 2022).
@@
 Practitioners see avoiding disciplinary action for non-compliance with company policies as a motivating benefit of documentation (Heger et al., 2022).
-Challenges.
-While practitioners acknowledge that the benefits of documentation could outweigh the costs of creating and maintaining it (Heger et al., 2022), when organizations fail to provide clear incentives for high-quality documentation, practitioners are less likely to prioritize it.
-This lack of motivation can hinder the realization of documentation’s potential benefits, but particularly its ability to inform downstream users.
-One major challenge is the lack of organizational incentives for documentation.
-Organizations typically do not prioritize documentation unless regulations require it for compliance or clients specifically request it (Chang and Custis, 2022; Miceli et al., 2021; Heger et al., 2022).
-Practitioners often view producing high-quality documentation as less relevant to their evaluations and promotions than development tasks that directly contribute to product outputs (Chang and Custis, 2022; Heger et al., 2022).
-This disconnect between the benefits of documentation and practitioner objectives is further exacerbated when the practitioners who benefit most from documentation (i.e., consumers) are not the same as those responsible for producing it (i.e., producers).
-In organizations where governance processes mandate documentation, a lack of familiarity with AI documentation concepts among developers could potentially magnify tensions between product development and governance teams (Ali et al., 2023; Ahlawat et al., 2024).
-Moreover, practitioners are often more motivated by compliance than by normative considerations, which could lead to blind spots when novel ethical issues arise outside existing policies (Heger et al., 2022).
-Facing time constraints and competing demands, practitioners may deprioritize documentation, sometimes cutting corners to save time (Chang and Custis, 2022; Heger et al., 2022).
-This can diminish the value of documentation artifacts, as evidenced by studies showing that practitioners sometimes reuse content from existing documentation, even when it pertains to different systems, leading to inadequate or incorrect documentation of their current system (Pushkarna et al., 2022; Heger et al., 2022).
-Practitioners also sometimes leave questions unanswered rather than seek out the necessary information (Heger et al., 2022) or prioritize recording information that is most valuable to them, potentially omitting critical details for others less familiar with the system (Hind et al., 2019).
-Furthermore, because practitioners often forget relevant information if they do not record it during development, their reconstruction of key system information after the fact can be time-consuming and error prone (Reid and Williams, 2023).
-Studies of publicly available documentation artifacts frequently highlight issues with incomplete or incorrect information, which can mislead downstream stakeholders regarding the characteristics of AI systems or components (Pepe et al., 2024; Liang et al., 2024; Geiger et al., 2020).
-For example, many publicly available model cards lack details about out-of-scope uses, limitations, or environmental impacts (Pepe et al., 2024; Liang et al., 2024).
-Less than 30% of model cards contain evaluation results, a critical piece of information for downstream users (Liang et al., 2024).
-Also, some publicly available documentation contains incorrect license information, potentially leading users to inadvertently violate license terms (Pepe et al., 2024).
-4.2.2. Potential impact #2: Documentation could support cross-functional collaboration and communication about AI risks.
-Documentation has the potential to act as a bridge between different stakeholder groups involved in the AI development lifecycle, including machine learning engineers, data scientists, product managers, user experience designers, researchers, and legal teams.
+The type of information that is most useful to downstream stakeholders may depend on whether the documentation pertains to data, models, systems, or methods and tasks.
+For example, system documentation can help downstream stakeholders understand functionality and risks that emerge from the interactions of underlying components (Arnold et al., 2019; Procope et al., 2022).
+Data documentation, on the other hand, can help downstream practitioners assess the downstream impact of upstream inputs, such as potentially privacy-compromising data inputs or sampling methods that could result in unfair user treatment (Gebru et al., 2021; Boyd, 2021).
+4.2.2. Theory of change #2: Documentation could support cross-functional collaboration and communication about AI risks.
 Since no single group oversees all stages of development, effective documentation can facilitate both indirect and direct communication among these diverse teams.
-Documentation artifacts can convey essential information across organizational boundaries, helping stakeholders who may not interact regularly to align their understanding of the system.
+Documentation artifacts can convey essential information across organizational boundaries, helping stakeholders who may not interact regularly to align their understanding of the system through documentation artifacts.
 Documentation can also prompt and support conversations among stakeholders directly, enabling them to work towards common goals despite their diverse backgrounds, frames of reference, or areas of expertise.
 This bridging function of documentation could break down organizational silos (Chmielinski et al., 2024) and promote effective collaboration across different teams (Crisan et al., 2022; Gilbert et al., 2023; Heger et al., 2022; Mohammad, 2022; Pushkarna et al., 2022; Raji and Yang, 2020; Richards et al., 2020; Shen et al., 2022; Srinivasan et al., 2021).
-Hypotheses.
-AI documentation could serve as an initial point of reference for downstream practitioners, offering a basic overview of the system that allows them to build foundational knowledge before engaging with developers for more in-depth discussions.
+At a basic level, AI documentation could serve as an initial point of reference for downstream practitioners, offering an overview of the system that allows them to build foundational knowledge before engaging with developers for more in-depth discussions (Liao et al., 2023; Heger et al., 2022).
 Rather than acting as a comprehensive, self-contained repository of information, documentation artifacts could provide an initial layer of understanding, enabling downstream practitioners to identify relevant areas for further exploration and formulate more informed questions when collaborating with developers.
@@
 Similarly, many AI practitioners review existing documentation only briefly before seeking further clarification through meetings or discussions (Heger et al., 2022).
-In more robust forms, documentation can facilitate deeper deliberation among stakeholders about the system’s design and the documentation process itself.
+In more robust forms, documentation can facilitate deeper deliberation among stakeholders about the system’s design and the documentation process itself (Pushkarna et al., 2022; Shen et al., 2022).
 For example, documentation has been shown to aid teams in collectively reflecting on socially constructed concepts, such as gender, and making informed decisions about how to annotate such attributes in datasets (Pushkarna et al., 2022).
@@
 Over time, engaging with documentation could also improve the technical literacy of stakeholders, enabling them to create products that better meet user needs while minimizing harm (Liao et al., 2023).
-Challenges.
-Several challenges complicate the role of documentation in facilitating cross-functional collaboration.
-The modern AI supply chain, characterized by the development of general-purpose models that can be adapted to a variety of applications downstream (Vaswani et al., 2017), often limits opportunities for direct communication between stakeholders.
-In some cases, different teams within the same organization may adapt a general-purpose model for various specific tasks.
-In these instances, documentation might prompt the downstream team to directly reach out to the original development team to discuss the system’s capabilities and limitations.
-In other cases, a team deploying a model may rely on a model developed by a different organization (Jones, 2023).
-In this case, it will typically be more difficult if not impossible for the downstream deployment team to directly communicate with the original development team.
-In either situation, the original developers may not fully anticipate the variety of downstream uses, leading to documentation that lacks critical details needed to identify potential risks or harms.
-Particularly when upstream and downstream teams cannot communicate directly, it will be more challenging for them to work together to understand the model’s risks within a given deployment context.
-This challenge is especially pronounced for less technical downstream stakeholders, who may struggle to understand the functionality and risks associated with these models.
-More interactive forms of documentation, such as those that allow users to engage with the models directly or customize the level of detail presented, could help mitigate these issues by making the information more accessible and relevant (Crisan et al., 2022; Liao et al., 2023; Moore et al., 2023), especially when direct communication between teams is limited.
-However, further research is needed to determine the most effective ways to document general-purpose systems while avoiding overwhelming documentation consumers with excessive or irrelevant details,
-Another major challenge is that documentation that is misaligned with the needs of its intended consumers can complicate rather than facilitate stakeholder communication.
-If documentation artifacts are overly technical or jargony, documentation consumers who are less technical or less familiar with the system may struggle to understand what details are most relevant, or misinterpret described practices (Crisan et al., 2022; Liao et al., 2023; Miceli et al., 2021; Shen et al., 2022); if documentation lacks relevant details, documentation consumers may overlook risks that become apparent only later in the development process, which may require teams to substantially backtrack.
-In either case, mismatch in comprehension can contribute to needless disruption to normal workflows.
-If practitioners perceive that documentation leads to seemingly egregious friction due to misunderstandings, they may be more prone to oversimplify the documentation they produce or gloss over relevant details that they worry will lead to confusion.
-4.2.3. Potential impact #3: Documentation could prompt practitioners to deliberate and act on the ethical impacts of AI
-Researchers have suggested that documentation has the potential to encourage practitioners to consider the ethical implications of their systems, leading to more responsible development and use of data, models, and systems.
-By integrating ethical reflection into the documentation process, practitioners may become more aware of the potential harms their systems could cause, prompting them to make more informed decisions.
-Hypotheses.
-One hypothesis is that when downstream documentation consumers consult documentation on data, models, and systems, it can prompt consideration of potential harms.
+4.2.3. Theory of change #3: Documentation could prompt practitioners to deliberate and act on the ethical impacts of AI
+As with other tools that surface AI harms (Balayn et al., 2023), documentation could encourage practitioners to consider the ethical implications of their systems.
+One theory is that when downstream documentation consumers consult documentation on data, models, and systems, it can prompt consideration of potential harms.
 This consideration could lead practitioners to make more careful decisions about if and when to use certain systems or components.
-Yet, empirical support for this hypothesis is inconsistent.
-One study of 23 practitioners examined whether reviewing data documentation artifacts could encourage practitioners to notice ethical issues, contextualize them, and formulate risk mitigation plans (Boyd, 2021).
-One subset received a scenario description, dataset, and data documentation, while the rest received only the scenario description and dataset.
-The researcher found that those with data documentation were more likely to notice ethical issues unprompted, though those without it often recognized concerns when the researcher pressed them — but practitioners had difficulty formulating action plans regardless of whether they had access to documentation.
-These findings suggest that while documentation artifacts can aid ethical deliberation, documentation is not always necessary and likely not sufficient to enable mitigation of ethical issues.
-Documentation also might sensitize producers to the ethical impacts of their work.
-Documentation frameworks often ask documentation producers to consider and describe the ethical dimensions of their systems.
+For example, one study found that practitioners with access to data documentation were more likely to identify ethical concerns unprompted.
+However, those without documentation often recognized similar concerns when prompted by the researcher.
+Despite this, practitioners both with and without access to documentation struggled to develop concrete action plans (Boyd, 2021).
+These findings indicate that while documentation can facilitate ethical deliberation, it is not always essential for ethical reflection and does not consistently lead to actionable solutions.
+Another theory is that documentation also might sensitize producers to the ethical impacts of their work because documentation frameworks often ask documentation producers to consider and describe the ethical dimensions of their systems.
 For example, the one framework prompts practitioners to think about privacy implications by inquiring about sensitive information in data (Gebru et al., 2021).
@@
 Their qualitative and survey study of Wikipedia communities found that framing ethical reflection and documentation as a participatory process led to more informed decisions about AI system design and deployment.
-Challenges.
-The assumption that documentation artifacts and processes can promote ethical reasoning among practitioners hinges on the belief that practitioners are sufficiently aware of AI’s ethical risks to users, non-users, and society that they can recognize them when prompted.
-However, without proper training in responsible AI practices and exposure to groups potentially harmed by AI systems, practitioners may not connect documentation with ethical risks (Heger et al., 2022).
-Consequently, their consideration of harms may only partially encompass the scope of risks, especially those requiring non-technical mitigations.
-For example, practitioners often misunderstand how bias can manifest in their work, leading them to incorrectly state in their documentation that bias concerns are not relevant (Hind et al., 2019).
-While it may not be necessary for ethicists to be embedded in the work process for practitioners to make responsible choices, as some have suggested (McLennan et al., 2022), findings point to the need for more support and training for practitioners in how to identify, document, and effectively communicate potential risks (Heger et al., 2022; Madaio et al., 2024).
-Furthermore, practitioners and organizations may resist documenting ethical impacts they have identifed.
-Publicly available documentation artifacts, such as Hugging Face model cards and GitHub repository README files, often show little consideration of ethical concerns (Pepe et al., 2024; Liang et al., 2024).
-While this omission might occur because practitioners are unaware of potential harms (Heger et al., 2022), their absence could also result from practitioners being hesitant to record ethical concerns they did recognize.
-Sometimes practitioners resist addressing ethical impacts because they feel unqualified to speculate on numerous potential use cases and their impacts (Crisan et al., 2022; Heger et al., 2022; Hind et al., 2019; Nunes et al., 2022).
-Others worry that detailing ethical concerns might give downstream stakeholders a false sense of security (Crisan et al., 2022).
-This hesitation could also stem from a desire to avoid personal responsibility for their actions (Nunes et al., 2022).
-Alternatively, practitioners could plausibly have concerns that documenting potential ethical harms could create a legal or public relations risk if outside stakeholders gained access to documentation.
-Even if documentation effectively engages practitioners in ethical reflection, they may have little influence over their organization’s development goals, deliverables, and timelines.
-As a result, they might be unable to make the changes to data, models, or systems that they have identified as useful or necessary to address ethical risks through the documentation process.
-One ethnographic study found that business demands, not the beliefs of data subjects or practitioners, largely determined the organization’s documentation approaches (Miceli et al., 2021).
-For example, although practitioners who reflected on data labeling practices recognized that socially constructed identities are complex, the organization nevertheless chose to represent identity in a reductive way, such as by defining race according to discrete, mutually exclusive categories.
-The authors concluded that explicit and implicit power structures among internal organizational stakeholders significantly affect practitioners’ documentation practices and ability to shape outcomes.
-Documentation approaches must be responsive to these constraints.
-Otherwise, documentation about ethical considerations may not promote meaningful action or may be incomplete or misleading.
-On the whole, evidence on the extent to which AI documentation frameworks can truly enhance ethical decision-making is mixed.
-Practitioners are capable of ethical deliberation about AI, but do not always thoroughly explore ethical issues or record the results in documentation artifacts.
-Even if practitioners engage in ethical deliberation when creating and using documentation, organizational environments can enable or constrain individual practitioners’ normative choices.
-4.2.4. Potential impact #4: Documentation could catalyze overall improvements in governance and development.
-Researchers have suggested that documentation can significantly influence governance and development practices broadly.
-In practical settings, AI documentation doesn’t function independently; documentation processes typically require practitioners to engage in other activities that can improve AI outcomes.
-Thus, the documentation process can catalyze behaviors that enhance the quality of AI system development and governance beyond documentation processes and artifacts.
-Hypotheses.
-Framework authors posit that documentation could improve the rigor with which practitioners develop AI systems (Holland et al., 2018).
+4.2.4. Theory of change #4: Documentation could catalyze overall improvements in governance and development.
+In applied settings, AI documentation rarely functions in isolation; it often relies on and reinforces complementary practices that could contribute to improved AI outcomes (Ahlawat et al., 2024).
+Our analysis revealed that several ways documentation could improve the rigor, efficiency, or reliability of development and governance practices overall.
+As one example, framework authors posit that documentation could improve the rigor with which practitioners develop AI systems (Holland et al., 2018).
 By requiring practitioners to justify their development choices, documentation may lead to more careful decision-making (Bender and Friedman, 2018; Chmielinski et al., 2024; Holland et al., 2018).
 Also, clear and comprehensive documentation supports reproducibility, aiding other practitioners in retraining and validating systems consistently (Adkins et al., 2022; Baracaldo et al., 2022; Hind et al., 2019).
-Since AI practitioners are often driven by a commitment to scientific rigor (Winecoff and Watkins, 2022), documentation that emphasizes this aspect can serve as a ”value lever (Shilton, 2013),” further encouraging engagement with the documentation process .
+Since AI practitioners are often driven by a commitment to scientific rigor (Winecoff and Watkins, 2022) and product quality assurance (Ahlawat et al., 2024), documentation that emphasizes these aspects could potentially serve as a ”value lever (Shilton, 2013),” further encouraging engagement with the documentation process.
 Documentation might also improve development and governance efficiency.
@@
 Practitioners also point out that by documenting unsuccessful and successful approaches, organizations can prevent repeated mistakes and identify practices that would be beneficial for the organization to disseminate broadly (Chang and Custis, 2022; Miceli et al., 2021).
-Documentation could also facilitate proactive risk mitigation, which is often more efficient than reactive approaches (Mitchell et al., 2019).
+Documentation could also facilitate proactive risk mitigation, which is often more efficient and effective than reactive approaches (Mitchell et al., 2019; Rakova et al., 2021).
 When issues arise in already-deployed systems, organizations typically need to fix the problem without disrupting service quality or availability.
 They might need to roll back to a previous version of the system that is less performant or apply quick patches that may not fully resolve the issue.
-Documentation that helps identify potential problems before deployment can lead to more comprehensive solutions than addressing issues post hoc once they are discovered in production systems (Holland et al., 2018).
-And helping teams discover and use approved AI components that have already undergone relevant risk management processes rather than creating new ones can help ensure risks are not re-introduced and overlooked.
+Documentation that helps identify potential problems before deployment can lead to more comprehensive solutions than addressing issues post hoc once they are discovered in production systems (Holland et al., 2018; Ahlawat et al., 2024).
 Furthermore, documentation could preserve institutional knowledge.
 In any organization, particularly complex ones, consistently institutionalizing values and best practices can be challenging.
-Documentation could serve as a means of conveying technical information about systems, and the organization’s policies and values that influence development and use.
+Documentation could serve as a means of conveying technical information about systems, and the organization’s policies and values that influence development and use (Ahlawat et al., 2024).
 For instance, documentation can aid in onboarding new employees by providing information about the systems they’ll work on and communicating the organization’s approach to development and governance (Adkins et al., 2022; Chang and Custis, 2022; Heger et al., 2022).
-Lastly, documentation could facilitate both internal and external audits (Brajovic et al., 2023; Crisan et al., 2022; Hupont and Gomez, 2022; Miceli et al., 2021).
+Lastly, documentation could facilitate both internal and external audits (Brajovic et al., 2023; Crisan et al., 2022; Hupont and Gomez, 2022; Miceli et al., 2021; Ahlawat et al., 2024).
 Audits may focus on the components of AI systems, the overall application, or the processes used in development (Mökander et al., 2023).
 Documentation could help organizations demonstrate that their claims—such as providing equitable system performance across different demographic groups—are backed by evidence.
-In these cases, documentation could guide auditors in selecting appropriate methods based on the auditor’s objectives and the available data.
-When audits examine the robustness of an organization’s governance, documentation becomes even more critical (Clavell, 2024).
+In these cases, documentation could guide auditors in selecting appropriate methods based on the auditor’s objectives and the available data (Ahlawat et al., 2024).
+When audits examine the robustness of an organization’s governance, documentation becomes even more critical since such information cannot be produced by analyzing the system itself (Clavell, 2024).
 Without comprehensive documentation, organizations may struggle to demonstrate the integrity and consistency of their processes.
-Challenges.
-One barrier to documentation facilitating best practices in development and governance is that the documentation process must be interconnected with other organizational processes to serve as an effective forcing function.
-In cases where documentation is ad-hoc and relies on tools that are not well-integrated with practitioners’ workflows (Bhat et al., 2023; Heger et al., 2022), connecting documentation processes and artifacts to other aspects of the development and governance process becomes even more challenging.
+4.3. Challenges
+Both proposed frameworks and empirical studies delineated how documentation could support more robust governance of AI systems; however, empirical studies, in particular, illuminated challenges organizations and practitioners often face, which could prevent the theories of change from being fully realized.
+These challenges included limited incentives and resources for robust documentation, structural and organizational barriers to communication, interpersonal and organizational constraints to ethical action, and poor integration with existing workflows, practices, and organizational infrastructure.
+4.3.1. Challenge #1: Poor incentivization and limited resources can undermine documentation quality.
+Practitioners acknowledge that the benefits of documentation could outweigh the costs of creating and maintaining it (Heger et al., 2022); however, when organizations fail to provide clear incentives for high-quality documentation, practitioners are less likely to prioritize it.
+This lack of motivation can hinder the realization of documentation’s potential benefits.
+Organizations typically do not prioritize documentation unless regulations require it for compliance or clients specifically request it (Chang and Custis, 2022; Miceli et al., 2021; Heger et al., 2022).
+Practitioners often view producing high-quality documentation as less relevant to their evaluations and promotions than development tasks that directly contribute to product outputs (Chang and Custis, 2022; Heger et al., 2022).
+This disconnect between the benefits of documentation and practitioner objectives is further exacerbated when the practitioners who benefit most from documentation (i.e., consumers) are not the same as those responsible for producing it (i.e., producers).
+In organizations where governance processes mandate documentation, a lack of familiarity with AI documentation concepts among developers could potentially magnify tensions between product development and governance teams (Ali et al., 2023; Ahlawat et al., 2024).
+Moreover, practitioners are often more motivated by compliance than by normative considerations, which could lead to blind spots when novel ethical issues arise outside existing policies (Heger et al., 2022).
+Under time constraints and competing priorities, practitioners may rush the documentation process, sometimes cutting corners to save time (Chang and Custis, 2022; Heger et al., 2022).
+For instance, they may reuse content from existing documentation, even when it describes different systems, resulting in incomplete or inaccurate records for the current system (Pushkarna et al., 2022; Heger et al., 2022; Ahlawat et al., 2024).
+Practitioners also sometimes leave questions unanswered rather than seek out the necessary information (Heger et al., 2022) or prioritize recording information that is most valuable to them, potentially omitting critical details for others less familiar with the system (Hind et al., 2019) or leaving downstream practitioners with an incomplete end-to-end understanding of systems (Ahlawat et al., 2024).
+Furthermore, because practitioners often forget relevant information if they do not record it during development, their reconstruction of key system information after the fact can be time-consuming and error prone (Reid and Williams, 2023).
+A potential consequence of rushed, low-quality documentation is that it may inadequately inform—or even misinform—downstream practitioners, potentially compromising effective development and risk management (Pepe et al., 2024; Liang et al., 2024; Geiger et al., 2020).
+For example, many publicly available model cards lack details about out-of-scope uses, limitations, or environmental impacts (Pepe et al., 2024; Liang et al., 2024).
+Less than 30% of model cards contain evaluation results, a critical piece of information for downstream practitioners (Liang et al., 2024).
+Also, some publicly available documentation contains incorrect license information, potentially leading practitioners to inadvertently violate license terms (Pepe et al., 2024).
+Practitioners frequently advocate for automating as much of the documentation process as possible to maintain its relevance and reduce its time demands (Rostamzadeh et al., 2022; Heger et al., 2022; Pushkarna et al., 2022; Ahlawat et al., 2024).
+However, certain essential details—such as the rationale behind specific design or development decisions—are challenging to automate.
+This type of information is often crucial for evaluating potential harms and for enabling accountability (Yurrita et al., 2023; Sambasivan et al., 2021).
+Recognizing this, some practitioners acknowledged that parts of the documentation process must remain manual (Pushkarna et al., 2022).
+However, if these manually updated sections are not regularly synchronized with system updates, outdated information could itself become a source of harm (McMillan-Major et al., 2021).
+As a result, deprioritizing documentation could lead to a situation where the most critical elements for risk management are the most likely to become outdated.
+4.3.2. Challenge #2: Structural and organizational barriers can complicate collaboration
+Several practical barriers can limit how well documentation can act as a springboard for active communication and collaboration.
+The modern AI supply chain, in which general-purpose models are adapted to a variety of downstream tasks (Vaswani et al., 2017), introduces additional structural complications for documentation.
+In some cases, different teams within the same organization may adapt a general-purpose model for specific tasks.
+In these instances, documentation might prompt the downstream team to directly reach out to the original development team to discuss the system’s capabilities and limitations or clarity details.
+However, in other cases, a team deploying a model may rely on a model developed by a different organization (Jones, 2023).
+In this case, it will typically be more difficult if not impossible for the downstream deployment team to directly communicate with the original development team.
+Any information that is missing or confusing in the documentation artifact cannot be easily clarified through collaboration.
+Furthermore, when AI development implicates complex supply chain dynamics, actors further removed from the deployment context often struggle to account for the needs and considerations of external stakeholders (Balayn et al., 2024).
+This disconnect may manifest in risk documentation for upstream models that fails to accurately or comprehensively capture the real-world risks these systems pose in specific deployment contexts.
+Challenges relevant to documentation of general-purpose models may be especially pronounced for less technical downstream stakeholders, who may struggle to understand the functionality and risks associated with these models (Crisan et al., 2022).
+More interactive forms of documentation, such as those that allow users to engage with the models directly or customize the level of detail presented, could help mitigate these issues by making the information more accessible and relevant (Crisan et al., 2022; Liao et al., 2023; Moore et al., 2023), especially when direct communication between teams is limited.
+However, further research is needed to determine the most effective ways to document general-purpose systems while avoiding overwhelming documentation consumers with excessive or irrelevant details.
+Beyond the structural challenges of general-purpose models, a major organizational challenge is that documentation that is misaligned with the needs of its intended consumers can complicate rather than facilitate stakeholder communication.
+If documentation artifacts are overly technical or jargony, documentation consumers who are less technical or less familiar with the system may struggle to understand what details are most relevant, or misinterpret described practices (Crisan et al., 2022; Liao et al., 2023; Miceli et al., 2021; Shen et al., 2022); if documentation lacks relevant details, documentation consumers may overlook risks that become apparent only later in the development process, which may require teams to substantially backtrack.
+In either case, mismatch in comprehension can contribute to needless disruption to normal workflows.
+4.3.3. Challenge #3: Interpersonal and organizational constraints can limit ethical decision making
+In some cases, documentation can improve critical reflection (Boyd, 2021); however, several practical barriers may limit the extent to which documentation can improve ethical deliberation and decision making.
+First, the assumption that documentation artifacts and processes can promote ethical reasoning among practitioners hinges on the belief that practitioners are sufficiently aware of AI’s ethical risks to users, non-users, and society that they can recognize them when prompted.
+However, without proper training in responsible AI practices and exposure to groups potentially harmed by AI systems, practitioners may not connect documentation with ethical risks (Heger et al., 2022).
+Consequently, their consideration of harms may only partially encompass the scope of risks, especially those requiring non-technical mitigations.
+For example, practitioners often misunderstand how bias can manifest in their work, leading them to incorrectly state in their documentation that bias concerns are not relevant (Hind et al., 2019).
+While it may not be necessary for ethicists to be embedded in the work process for practitioners to make responsible choices, as some have suggested (McLennan et al., 2022), findings point to the need for more support and training for practitioners in how to identify, document, and effectively communicate potential risks (Heger et al., 2022; Madaio et al., 2024).
+Second, practitioners and organizations may resist documenting ethical impacts they have identified.
+Publicly available documentation artifacts, such as Hugging Face model cards and GitHub repository README files, often show little consideration of ethical concerns (Pepe et al., 2024; Liang et al., 2024).
+Sometimes practitioners resist addressing ethical impacts because they feel unqualified to speculate on numerous potential use cases and their impacts (Crisan et al., 2022; Heger et al., 2022; Hind et al., 2019; Nunes et al., 2022; Ahlawat et al., 2024).
+Others worry that detailing ethical concerns might give downstream stakeholders a false sense of security (Crisan et al., 2022).
+This hesitation could also stem from a desire to avoid personal responsibility for their actions (Nunes et al., 2022).
+Alternatively, practitioners could plausibly have concerns that documenting potential ethical harms could create a legal or public relations risk if outside stakeholders gained access to documentation.
+Third, even if practitioners recognize and record ethical concerns, they may have little influence over their organization’s development goals, deliverables, and timelines (Ahlawat et al., 2024).
+As a result, they might be unable to make the changes to data, models, or systems that they have identified as useful or necessary to address ethical risks through the documentation process (Ahlawat et al., 2024; Miceli et al., 2021).
+One ethnographic study found that business demands, not the beliefs of data subjects or practitioners, largely determined the organization’s documentation approaches (Miceli et al., 2021).
+For example, although practitioners who reflected on data labeling practices recognized that socially constructed identities are complex, the organization nevertheless chose to represent identity in a reductive way, such as by defining race according to discrete, mutually exclusive categories.
+The authors concluded that explicit and implicit power structures among internal organizational stakeholders significantly affect practitioners’ documentation practices and ability to shape outcomes.
+Documentation approaches must be responsive to these constraints.
+Otherwise, documentation about ethical considerations may not promote meaningful action or may be incomplete or misleading.
+4.3.4. Challenge #4: Poor integration with technical and organizational infrastructure can undermine diffuse benefits
+Practitioners studies emphasize that documentation is likely to be more effective when it is well integrated with practitioner workflows (Heger et al., 2022; Bhat et al., 2023).
+For example, when practitioners are asked to produce documentation using external tools, they are more likely to make errors and are less likely to make note of relevant ethical considerations than when documentation methods are well integrated with their existing tools (Bhat et al., 2023).
+Beyond undermining the quality of documentation artifacts themselves, poor technical and organizational integration of documentation processes with other organizational infrastructure could limit the extent to which documentation serves as a forcing function for best practices in development and governance generally.
+In cases where documentation is ad-hoc and relies on tools that are not well-integrated with practitioners’ workflows (Bhat et al., 2023; Heger et al., 2022), connecting documentation processes and artifacts to other aspects of the development and governance process becomes even more challenging (Ahlawat et al., 2024).
 Without a solid documentation infrastructure that aligns with other organizational functions, documentation may fail to effectively enhance development and risk management.
-Another challenge is that while high-quality documentation can improve governance and development, poor-quality documentation can have negative impacts.
+Another challenge is that while high-quality documentation can have compounding benefits to governance and development, poor-quality documentation can have compounding negative impacts.
 Documentation should prevent redundant efforts, maximize time spent managing unique risks, and support proactive system development.
@@
 Inadequate or incorrect documentation can lead to wasted time, overlooked risks, and unexpected issues that may require correction later in development.
-This could result in harms remaining unaddressed or inefficiencies that undermine the overall governance process.
+This could result in harms remaining unaddressed or inefficiencies undermining the overall governance process or quality and reliability of development.
 4.4. Design considerations
@@
 To an extent, these design considerations are interrelated: the level of detail documentation artifacts contain can also be a form of audience tailoring, for example.
-Yet each of these design tensions has partially distinct motivations and implications, so we address them separately.
+And, audience tailoring can be viewed as a form of customization.
+Yet each of these design tensions at least partially distinct motivations and implications, so we address them separately.
 4.4.1. Design consideration #1: Determining the degree of customization versus standardization in documentation artifacts.
@@
 At the organizational level, customization might also be necessary.
-For example, a healthcare startup may require a different documentation approach than a large, well-established financial institution.
+For example, a healthcare company may require a different documentation approach than a financial company, and the approach either should take may depend on the company’s maturity level (Dotan et al., 2024).
 In some cases, it may be appropriate to customize documentation at multiple levels to respond to both the organization’s nature and the type of system it is developing (Chmielinski et al., 2020; Gebru et al., 2021; Holland et al., 2018; Richards et al., 2020; Roman et al., 2023; Stoyanovich and Howe, 2019).
@@
 Practitioners often struggle with how to complete documentation or what level of detail to include (Chang and Custis, 2022; Pushkarna et al., 2022; Heger et al., 2022; Miceli et al., 2021), leading them to rely heavily on existing examples, which may not always be appropriate.
-In one study, practitioners frequently copied documentation from previous projects into their current work, sometimes including irrelevant information or omitting crucial details (Pushkarna et al., 2022).
-Standardization can mitigate these issues by providing clear expectations for what information should be included and what users should expect to find in documentation artifacts, thus encouraging more consistent practices (McMillan-Major et al., 2021; Pushkarna et al., 2022; Roman et al., 2023).
+Standardization can mitigate these issues by providing clear expectations for what information should be included and what consumers should expect to find in documentation artifacts, thus encouraging more consistent practices (McMillan-Major et al., 2021; Pushkarna et al., 2022; Roman et al., 2023).
 Standardization also might facilitate structured risk management.
@@
 The choice between customization and standardization is not always a strict binary but a spectrum, allowing for a combination of both.
-For example, organizations can create standardized templates that are adapted to different contexts, such as healthcare models versus those for software engineers, ensuring consistency while meeting specific needs.
-However, organizations must thoughtfully decide when to standardize and when to customize.
+For example, organizations can create standardized templates that are adapted to different contexts, such as healthcare models versus those for software engineers, ensuring consistency while meeting specific needs (Bhat et al., 2023; Heger et al., 2022; Hind et al., 2019; Holland et al., 2018; McMillan-Major et al., 2021; Pushkarna et al., 2022; Richards et al., 2020).
+However, organizations may face trade-offs when deciding on the level of standardization to employ in documentation.
 4.4.2. Design consideration #2: Determining how much to tailor documentation artifacts to specific audiences
-When designing documentation strategies, organizations must consider the diverse needs of multiple stakeholder groups.
+Organizations could tailor their documentation to fit specific application contexts or systems.
+In addition, tailoring documentation to meet the needs of its intended audience can help ensure it achieves the desired governance outcomes.
+When designing documentation strategies, organizations must consider the diverse needs of multiple stakeholder groups (Micheli et al., 2023).
 Different stakeholders often require distinct forms of documentation to be effective.
@@
 First, when documentation is aligned with the purpose of their role, practitioners are better able to understand the necessary actions to manage risk.
-For example, information about how well a system meets user needs may be more actionable for a UX professional than a machine learning engineer.
+For example, information about how well a system meets user needs may be more actionable for a UX professional than a machine learning engineer (Liang et al., 2024).
 Second, practitioners typically have the authority to act in specific ways within organizations.
@@
 In cases where time constraints are a concern, it may be more effective to focus on producing a single, high-quality document that is regularly updated.
-At a minimum, all documentation should clearly indicate its last update to help users gauge its relevance and accuracy.
+At a minimum, it may be beneficial if documentation artifacts clearly indicate their last update to help practitioners gauge relevance and accuracy.
 4.4.3. Design trade-off #3: Determining the level of interactivity documentation should support
-Producing multiple forms of documentation is one way to support a variety of goals, stakeholders, and systems and goals; however, these multiple forms of documentation create challenges for organizations to manage.
+Producing multiple forms of documentation is one way to support a variety of goals, stakeholders, and systems; however, these multiple forms of documentation create challenges for organizations to manage.
 Interactive documentation offers an alternative solution by enabling stakeholders to access the specific information they need while maintaining overall usability and comprehension.
@@
 Empirical research supports the idea that interactive documentation can significantly improve understanding, particularly among non-technical audiences (Crisan et al., 2022; Liao et al., 2023; Moore et al., 2023).
-Tools like Hugging Face’s model inference API444https://huggingface.co/docs/api-inference/ or OpenAI’s developer playground555https://platform.openai.com/playground allow practitioners to engage with models and systems directly, fostering an intuitive understanding of how inputs and outputs are connected.
+Tools like Hugging Face’s model inference API555https://huggingface.co/docs/api-inference/ or OpenAI’s developer playground666https://platform.openai.com/playground allow practitioners to engage with models and systems directly, fostering an intuitive understanding of how inputs and outputs are connected.
 This hands-on interaction is crucial for designing products that leverage the system’s capabilities and for assessing potential risks (Liao et al., 2023).
-Interactive access is particularly valuable when exploring pre-trained, general-purpose models, whose functionality can vary significantly depending on the implementation (Liao et al., 2023; Moore et al., 2023; Crisan et al., 2022).
+Interactive access is particularly valuable when exploring pre-trained, general-purpose models, where functionality can vary significantly depending on the implementation (Liao et al., 2023; Moore et al., 2023; Crisan et al., 2022).
 However, interactive documentation also presents challenges.
@@
 Whereas exhaustive documentation might ensure practitioners have access to all the necessary information, concise documentation may improve ease of production and use.
-Concise documentation can help stakeholders focus on the details that are most relevant to their needs, reducing the likelihood of information overload.
+Concise documentation could help stakeholders focus on the details that are most relevant to their needs, reducing the likelihood of information overload.
 For documentation aimed at downstream consumers, including internal risk management professionals, it is essential to provide enough detail to support decision-making without overwhelming them.
 Information overload can lead to selective attention, where decision-makers may focus on certain details while neglecting others, potentially degrading the quality of their decisions (Phillips-Wren and Adya, 2020), particularly under time constraints (Hahn et al., 1992).
-For example, faced with extensive and complex documentation on AI systems, consumers might miss critical information or become frustrated and abandon the effort.
+For example, faced with extensive and complex documentation on AI systems, consumers might not read documentation at all, miss critical information, or become frustrated and abandon the effort.
 As one participant in a research study noted, ”I would lose patience after 30-40 seconds if I have to put a lot of effort into finding what I’m looking for” (Crisan et al., 2022).
 Concise documentation that highlights the most important information for a given stakeholder group can help internal decision-makers manage risk more effectively by ensuring they focus on critical details.
-Moreover, concise documentation is often easier to produce and maintain, which can enhance accuracy and encourage more frequent use.
+Moreover, concise documentation is likely easier to produce and maintain, which can enhance accuracy and encourage more frequent use.
 Exhaustive documentation may be burdensome to create and keep updated, particularly if the process is manual.
@@
 This includes the motivations for developing specific datasets or models, criteria for selecting data sources, reasoning behind chosen methodologies, outcomes of compliance reviews, contact information for responsible parties, decommissioning procedures, and ethical considerations addressed during development.
+Providing detailed documentation of the rationale for design and development decisions underscores the subjective nature of these choices, which are often the underlying causes of AI-related harms (Yurrita et al., 2023; Sambasivan et al., 2021).
+Manually created documentation can empower downstream stakeholders—and ultimately users—to scrutinize and contest the treatment they receive from AI systems (Yurrita et al., 2023).
+By transparently detailing how normative assumptions and subjective decisions are incorporated into system design and development, organizations can enhance accountability, supporting both internal governance and external stakeholder oversight.
 One of the purposes of both producing and using documentation within an organization is to prompt developers to consider the ethical implications of their systems.
@@
 Well-designed automation of some portions of documentation could reduce this risk.
-4.4. Recommendations
+5. Limitations
+Our research design and analysis have several limitations.
+First, we did not directly study the implementation of any specific documentation framework.
+Consequently, our assessment of the utility of our theories of change, and the challenges associated with them, is based on indirect analysis of existing work rather than direct evaluation.
+Another limitation is the broad scope we applied to evaluating documentation.
+Different domains of documentation, such as documentation for voice data versus art data, documentation for specific system components, such as data versus models, and documentation for different stages of development, such as pretraining data versus model fine-tuning data, may have distinct impacts on governance.
+Moreover, as with other aspects of governance (Rakova et al., 2021; Winecoff and Watkins, 2022), the role of documentation likely varies across industries and may function differently depending on an organization’s size and level of governance maturity (Ahlawat et al., 2024; Rakova et al., 2021).
+Our broad approach limits the depth of our analysis in addressing these specific nuances.
+Lastly, our work did not explore how multiple forms of transparency could support or improve AI governance.
+Documentation is one of many tools that can provide transparency into AI systems and support internal and external accountability.
+In some cases, system- or component-level documentation alone may not sufficiently enable practitioners to understand how systems operate, the risks they pose, or why they might behave in unexpected or potentially harmful ways.
+In these cases, explainability methods or other forms of transparency may be necessary to provide the insight needed to understand risks or resolve outstanding issues.
+6. Directions for future research
 Given that organizations have unique requirements and operate in vastly different contexts, there is no one-size-fits-all approach to documentation or a universal set of information that applies to all datasets, models, or systems.
-However, our analysis of proposed frameworks and empirical findings suggests several considerations that organizations should take into account when designing and implementing their documentation strategies.
-Each of these recommendations could be further strengthened by novel empirical research about how organizations can maximize the benefits of implementing the recommendations in applied contexts.
-Identify critical stakeholders and their documentation needs.
-When developing AI documentation strategies, organizations should begin by identifying the stakeholders who will frequently and significantly interact with both the documentation processes and artifacts.
-It’s important to consider not only the needs of those who will consume the documentation but also the expertise and capacity of those responsible for producing it.
-If the documentation is designed to be highly informative for downstream users but is impractical for upstream practitioners to produce, the approach is likely to fail in practice.
-In such cases, organizations may need to expand their documentation teams to ensure they can meet the needs of consumers effectively.
-Also, organizations should take into account the technical and non-technical backgrounds of their stakeholders to strike a balance between providing the necessary level of detail and maintaining accessibility.
-Evaluate operational resources and constraints and incentivize the production of robust documentation within that context.
-A realistic assessment of available resources—such as time, personnel, and expertise—is crucial for supporting effective documentation.
-Overly ambitious documentation efforts without adequate resources can lead to failures, such as incomplete or inaccurate documentation.
-It is often better to produce concise, high-quality documentation than to attempt a comprehensive approach that falls short.
-Organizations should ensure that practitioners understand their documentation responsibilities and that incentives are in place to promote thorough and accurate documentation practices.
-Leverage empirical evidence and feedback for continuous improvement.
-Organizations should evaluate potential documentation approaches via pilot projects and periodically collect empirical data on the efficacy of their documentation practices in meeting the needs of relevant stakeholders.
-Organizations should plan to assess the usability of documentation frameworks and their effectiveness in achieving the organization’s goals (Berman et al., 2024), and adapt their documentation strategies by conducting continuous assessments as their needs or technology evolves.
-Given the limited empirical data on the effectiveness of documentation approaches, organizations can also contribute to the broader AI community by publicly sharing their findings on what works best in different contexts.
-Organizations should recognize that empirical evidence and key performance indicators can be qualitative and quantitative.
-Whereas quantitative evidence can help organizations assess general trends at scale, qualitative evidence can help them understand unique issues in greater depth.
-Use standardization as the default approach, and intentionally deviate from it when necessary.
-Standardization offers many advantages.
-It allows practitioners to compare documentation artifacts, facilitates search and discoverability of information, establishes consistent expectations for documentation producers and consumers, and can help institutionalize norms of practice.
+Our analysis of proposed frameworks and empirical findings suggests several potential pathways organizations could explore to develop documentation practices tailored to their specific needs and constraints.
+Further design and empirical research in key areas could address the limitations of our study or expand into other promising directions.
+Such work could provide organizations across varying maturity levels and domains with more specific, actionable guidance for designing documentation processes that strengthen their AI governance.
+Leveraging empirical evidence and feedback for continuous improvement.
+Our sample of publications contained significantly higher number of proposed documentation frameworks compared to empirical evaluations of their effectiveness, highlighting gaps in understanding how theoretical approaches to documentation can be translated into practical, impactful practices within specific organizational contexts.
+Notably, most empirical findings on documentation were cross-sectional, offering limited insight into how iterative evaluation and refinement could enhance documentation practices over time within AI organizations.
+Future empirical research could fill this gap by identifying how organizations can evaluate potential documentation approaches via pilot projects and periodically collect empirical data on the efficacy of their documentation practices in meeting the needs of relevant stakeholders.
+Research could also inform how organizations can plan to assess the usability of documentation frameworks and their effectiveness in achieving the organization’s goals (Berman et al., 2024), and adapt their documentation strategies by conducting continuous assessments as their needs or technology evolves.
+Beyond scholarly research efforts, given the limited empirical data on the effectiveness of documentation approaches, organizations can also contribute to the broader AI community by publicly sharing their findings on what works best in different contexts.
+Organizations may benefit from empirical evidence and key performance indicators that are both qualitative and quantitative.
+Quantitative metrics are often best suited to capturing broad patterns across many instances but may lack depth for individual phenomena.
+Qualitative measures provide ”thick” insights for specific instances (Geertz, 2008) but may have limited generalizability.
+Identifying unexpected negative consequences of design choices.
+Empirical findings from researchers or organizations on how theoretically sound design choices—including those discussed in this work—might lead to unintended negative consequences can offer valuable insights for shaping best practices grounded in real-world experiences.
+For instance, a study on practitioners might reveal that while interactive interfaces help non-technical users understand information, they could hinder technical practitioners from quickly accessing relevant details.
+Similarly, research on practitioners or publicly available artifacts might show that practitioners often omit critical information, even when documentation is designed to be concise.
+Such findings could inform revisions to the frameworks and design considerations discussed in this work.
+Informing the degree and design of standardization of artifacts.
+Our research highlights the advantages of standardization including enabling comparison between documentation artifacts, facilitating search and discoverability of information, establishing consistent expectations for documentation producers and consumers, and institutionalizing norms of practice.
 For these reasons, some level of standardization in documentation processes and artifacts is likely to be helpful for most organizations for most AI systems.
-To balance standardization and customization, organizations might develop customizable documentation templates or interchangeable modules (Bhat et al., 2023; Heger et al., 2022; Hind et al., 2019; Holland et al., 2018; McMillan-Major et al., 2021; Pushkarna et al., 2022; Richards et al., 2020).
+However, standardized formats may overlook or insufficiently address the unique characteristics of specific AI system components or categories (e.g., language versus vision models), or application domains (e.g., social media versus healthcare).
+Future research could benefit more specific applications by providing insight into the appropriate level of standardization and identifying when and to what extent deviations from standardized formats are advantageous.
+One approach to balancing standardization and customization that research has begun to explore is having customizable documentation templates or interchangeable modules (Bhat et al., 2023; Heger et al., 2022; Hind et al., 2019; Holland et al., 2018; McMillan-Major et al., 2021; Pushkarna et al., 2022; Richards et al., 2020).
 Templates could, for example, allow for customization of subsections while maintaining consistent content and structure within those subsections.
 For instance, companies may include a subsection on the labeling process for AI systems based on supervised learning but exclude it for unsupervised AI systems since they do not require labels.
-Where companies choose to embrace a greater degree of customization, they should ensure that the benefits of this customization outweigh the benefits of maintaining standardization.
-Leverage automation for information extraction and incorporate manual processes for scoped critical synthesis.
-Some helpful information for documentation, such as dataset composition, statistical properties, model training methods, parameters, and software library or third-party API usage, can be extracted from AI systems using automated procedures, contributing to consistent, high-fidelity documentation of these system elements.
-While automated documentation can provide reliable and up-to-date information, it has limitations.
-The aspects of systems that are easiest to document automatically are not always the most critical, and errors in automated documentation can have a greater negative impact than manual errors, as they can persist through each update.
-Therefore, organizations should be selective about the information included in documentation artifacts, even if it can be easily generated automatically.
-Consider interactive documentation interfaces for general-purpose models in particular.
-Documenting the characteristics of general-purpose models and their potential risks is particularly challenging due to their wide range of possible applications.
-When downstream practitioners cannot directly communicate with upstream model developers, interpreting the general documentation in the context of specific applications can be difficult.
+Future research could explore how templates or other customization approaches can best maintain the usability and institutional benefits of standardization, while allowing for enough customization to address unique risks.
+Designing documentation for general-purpose AI systems.
+Although initiatives such as the EU AI Act require general-purpose model providers to supply downstream deployers with certain technical documentation to support their risk management, research suggests that the properties of upstream or base models may not always be relevant to downstream deployments (Weidinger et al., 2023; Solaiman et al., 2023; Qi et al., 2023).
+Although approaches for documentation general-purpose models are emerging (Bommasani et al., 2023), disconnects between upstream models and downstream applications pose significant challenges for designing effective documentation recommendations for governing general-purpose systems.
+For example, when downstream practitioners cannot directly communicate with upstream model developers, interpreting the general documentation in the context of specific applications can be difficult.
 Several studies have shown that interfaces allowing practitioners to observe system outputs in response to given inputs can enhance their understanding of model capabilities and risks (Liao et al., 2023; Crisan et al., 2022; Moore et al., 2023).
 Consequently, interactive interfaces can serve as a valuable complement to traditional, static documentation.
-Focus initially on core information needs for documentation and evaluate potential gaps over time.
-When designing and implementing new documentation strategies, organizations should start by identifying the essential pieces of information that will meet the needs of key stakeholders.
-In other words, they should start by designing a concise version of documentation.
-They should plan to consult with stakeholders who use this documentation to determine whether any additional information is necessary to perform their essential duties and whether any of the initial information provided in documentation artifacts is unnecessary.
-They should also incorporate learnings from empirical research to augment documentation as needed.
-Organizations can use this agile approach to manage the information in documentation artifacts, ensuring that the time spent producing and using documentation is maximally efficient and effective at serving stakeholder requirements.
-Institutionalize responsible AI and risk management practices beyond documentation.
+Research that illuminates how traditional, static approaches to documentation can be combined with interactive interfaces and even explainability methods may shed light on how documentation of general-purpose systems can best be communicated throughout the AI development lifecycle.
+Such work may be especially helpful in designing information architectures for documentation for practitioners without deep AI expertise, since these practitioners often face challenges in understanding complex technical information in model documentation (Crisan et al., 2022).
+Disentangling the role of different types of documentation in AI governance.
+Our analysis reveals that AI documentation can broadly support a range of governance objectives through four general mechanisms.
+However, different types of documentation—such as data, model, system, or process documentation—either individually or in specific combinations, may serve distinct functions.
+For example, data documentation provides practitioners with insights into how inputs might influence system behavior.
+Details about the dataset—such as who is represented and what information is contained within data instances—can alert downstream practitioners to potential biases or risks, including the possibility of unfair outcomes or privacy violations.
+Model documentation, on the other hand, can help practitioners evaluate the reliability of system performance or make informed decisions about usage, considering factors like environmental impacts, which can be significant even during model inference (Chien et al., 2023).
+System and process documentation plays a crucial role in providing operational clarity, enabling downstream stakeholders or organizational decision-makers to assess deployment contexts, overarching system behaviors, performance, risks, and the adequacy of ongoing monitoring processes.
+To ensure maximum effectiveness, the form and content of documentation artifacts—as well as the processes involved in creating them—should be responsive to the unique impacts and governance challenges associated with different system components.
+Future research, particularly empirical studies focused on organizational contexts, could clarify how documenting different system components—individually or in combination—supports distinct or overlapping governance goals.
+Such analysis could also inform design recommendations for different types of documentation, ensuring they collectively maximize their overall impact on governance.
+Institutionalizing responsible AI and risk management practices beyond documentation.
 Ethical deliberation is a skill that practitioners can develop over time with intentional practice (Vallor et al., 2018), but organizations need to implement additional interventions beyond documentation to help cultivate this skill.
 Even when practitioners engage in documentation processes, they may overlook critical risks or fail to recommend appropriate mitigations without a deeper understanding of how AI can harm diverse individuals and communities (Madaio et al., 2024; Chang and Custis, 2022).
-Therefore, organizations should strive to embed responsible AI and risk management efforts throughout their AI development practices.
+Therefore, organizations should strive to embed responsible AI and risk management efforts throughout their tools and AI development practices.
 When responsible AI is integrated into the organization’s culture and broader practices, documentation processes that prompt practitioners to reflect on risks are more likely to be informed by a comprehensive understanding necessary for effective risk management.
-Conversely, relying solely on documentation to encourage critical reflection is unlikely to be sufficient and may even lead to negative outcomes.
-Documentation of potential ethical harms will likely be more effective when it focuses on concrete system properties, such as data containing personally identifiable information or existing societal biases, rather than on speculative concerns, which can be subjective and heavily influenced by the expertise and perspectives of the documentation producers.
-4.5. Conclusion
+An open research question is how different tools and methods for raising awareness of ethical issues, fostering ethical deliberation, and empowering practitioners to act on these insights can work together to enhance AI governance.
+For instance, fairness toolkits can help practitioners less-experienced in algorithmic fairness better recognize and deliberate on trade-offs between system fairness and other performance indicators (Balayn et al., 2023).
+However, the tools and processes organizations adopt to institutionalize risk management practices may inadvertently narrow practitioners’ focus, limiting their ability to address emergent problems that fall outside the scope of these tools and processes or encouraging them to apply risk management techniques in inappropriate contexts (Balayn et al., 2023; Ahlawat et al., 2024).
+Future research could explore how documentation and other AI risk mitigation tools can enable practitioners to manage risks consistently and rigorously while remaining flexible enough to identify and address new and unforeseen concerns.
+Research on how explainability methods can support documentation may be especially helpful in this regard.
+Stakeholders seek explanations for a variety of reasons, such as translating system predictions into concrete actions, ensuring compliance with relevant policies and regulations, understanding the impact of different model factors on their outputs, among others (Suresh et al., 2021).
+While documentation could alert practitioners to possible issues that might arise in deployment, it does not necessarily enable causal understanding of issues.
+Explainability methods, in contrast, can support practitioners in understanding model behavior at the instance level, helping them debug models or systems (Bhatt et al., 2020).
+Explainability methods may be especially useful in for helping practitioners conceptualize the behavior of general-purpose AI systems such as large language models (LLMs) (Liao and Vaughan, 2024).
+By identifying which stakeholders would benefit most from explanations and what their goals are, organizations can design explainability approaches that work in tandem with documentation to support responsible risk management and decision making.
+7. Conclusion
 This paper underscores the critical role of documentation in enhancing AI governance within organizations, emphasizing that effective documentation practices are essential for managing AI risks and fostering responsible system development.
@@
 Minimum Viable Ethics: From Institutionalizing Industry AI Governance to Product Impact.
-arXiv (September 11 2024).
-https://www.arxiv.org/abs/2409.06926
+arXiv preprint arXiv:2409.06926 (2024).
+https://arxiv.org/html/2409.06926v1
 - Ali et al.
@@
 http://arxiv.org/abs/1808.07261
+- Balayn et al.
+(2024) Agathe Balayn, Lorenzo Corti, Fanny Rancourt, Fabio Casati, and Ujwal Gadiraju.
+2024.
+Understanding Stakeholders’ Perceptions and Needs Across the LLM Supply Chain.
+arXiv preprint arXiv:2405.16311 (2024).
+https://arxiv.org/abs/2405.16311
+- Balayn et al.
+(2023) Agathe Balayn, Mireia Yurrita, Jie Yang, and Ujwal Gadiraju.
+2023.
+”Fairness Toolkits, A Checkbox Culture?” On the Factors that Fragment Developer Practices in Handling Algorithmic Harms.
+In Proceedings of the 2023 AAAI/ACM Conference on AI, Ethics, and Society.
+482–495.
+https://dl.acm.org/doi/10.1145/3600211.3604674
 - Baracaldo et al.
@@
 https://doi.org/10.1145/3544548.3581518
-- Blasch et al. (2021) Erik Blasch, James Sung, and Tao Nguyen. 2021. Multisource AI scorecard table for system evaluation. In AAAI FSS20: Artificial Intelligence in Government and Public Sector.
+- Bhatt et al.
+(2020) Umang Bhatt, Alice Xiang, Shubham Sharma, Adrian Weller, Ankur Taly, Yunhan Jia, Joydeep Ghosh, Ruchir Puri, José MF Moura, and Peter Eckersley.
+2020.
+Explainable Machine Learning in Deployment.
+In Proceedings of the 2020 Conference on Fairness, Accountability, and Transparency.
+648–657.
+https://arxiv.org/abs/1909.06342
+- Blasch et al.
+(2021) Erik Blasch, James Sung, and Tao Nguyen.
+2021.
+Multisource AI Scorecard Table for System Evaluation.
+In AAAI FSS20: Artificial Intelligence in Government and Public Sector.
+https://arxiv.org/abs/2102.03985
+- Bommasani et al.
+(2023) Rishi Bommasani, Kevin Klyman, Shayne Longpre, Sayash Kapoor, Nestor Maslej, Betty Xiong, Daniel Zhang, and Percy Liang.
+2023.
+The foundation model transparency index.
+arXiv preprint arXiv:2310.12941 (2023).
+https://arxiv.org/abs/2310.12941
 - Boyd (2021) Karen L.
@@
 Understanding Implementation Challenges in Machine Learning Documentation.
-In Equity and Access in Algorithms, Mechanisms, and Optimization.
+In Proceedings of the 2nd ACM Conference on Equity and Access in Algorithms, Mechanisms, and Optimization.
 ACM, Arlington VA USA, 1–8.
@@
 - Charmaz (2014) Kathy Charmaz. 2014. Constructing Grounded Theory, 2nd Ed. Sage.
+- Chien et al.
+(2023) Andrew A Chien, Liuzixuan Lin, Hai Nguyen, Varsha Rao, Tristan Sharma, and Rajini Wijayawardana.
+2023.
+Reducing the Carbon Impact of Generative AI Inference (today and in 2035).
+In Proceedings of the 2nd Workshop on Sustainable Computer Systems.
+1–7.
+https://dl.acm.org/doi/10.1145/3604930.3605705
 - Chmielinski et al.
@@
 The Dataset Nutrition Label (2nd Gen): Leveraging Context to Mitigate Harms in Artificial Intelligence.
+https://arxiv.org/abs/2201.03954
 - Clavell (2024) Gemma Galdon Clavell. 2024. Checklist for AI Auditing. https://www.edpb.europa.eu/system/files/2024-06/ai-auditing_checklist-for-ai-auditing-scores_edpb-spe-programme_en.pdf
@@
 http://arxiv.org/abs/2311.17259
+- Dotan et al.
+(2024) Ravit Dotan, Borhane Blili-Hamelin, Ravi Madhavan, Jeanna Matthews, and Joshua Scarpino.
+2024.
+Evolving AI Risk Management: A Maturity Model based on the NIST AI Risk Management Framework.
+arXiv preprint arXiv:2401.15229 (2024).
+https://arxiv.org/abs/2401.15229
 - Friedman et al.
@@
 http://arxiv.org/abs/1803.09010
+- Geertz (2008) Clifford Geertz. 2008. Thick Description: Toward an Interpretive Theory of Culture. In The Cultural Geography Reader. Routledge, 41–51.
 - Geiger et al.
@@
 325–336.
+https://arxiv.org/abs/1912.08320
 - Gilbert et al.
@@
 http://arxiv.org/abs/2204.10817
-- Green et al.
-(2021) Nekesha Green, Chavez Procope, Adeel Cheema, and Adekunle Adediji.
-2021.
-System Cards, a New Resource for Understanding How AI Systems Work.
-https://ai.facebook.com/blog/system-cards-a-new-resource-for-understanding-how-ai-systems-work/
+- Glaser (1978) Barney G Glaser. 1978. Theoretical sensitivity. University of California,.
+- Glaser et al. (1968) Barney G Glaser, Anselm L Strauss, and Elizabeth Strutzel. 1968. The Discovery of Grounded Theory: Strategies for Qualitative Research. Nursing Research 17, 4 (1968), 364.
 - Hahn et al. (1992) Minhi Hahn, Robert Lawson, and Young Gyu Lee. 1992. The Effects of Time Pressure and Information Load on Decision Quality. Psychology & Marketing 9, 5 (1992), 365–378.
@@
 http://arxiv.org/abs/1911.08293
-- Holland et al. (2018) Sarah Holland, Ahmed Hosny, Sarah Newman, Joshua Joseph, and Kasia Chmielinski. 2018. The Dataset Nutrition Label: A Framework To Drive Higher Data Quality Standards.
+- Holland et al.
+(2018) Sarah Holland, Ahmed Hosny, Sarah Newman, Joshua Joseph, and Kasia Chmielinski.
+2018.
+The Dataset Nutrition Label: A Framework To Drive Higher Data Quality Standards.
+https://arxiv.org/abs/1805.03677
 - Hupont and Gomez (2022) Isabelle Hupont and Emilia Gomez.
@@
 http://arxiv.org/abs/2302.10395
-- Liao and Vaughan (2023) Q Vera Liao and Jennifer Wortman Vaughan. 2023. Ai transparency in the age of llms: A human-centered research roadmap. arXiv preprint arXiv:2306.01941 (2023), 5368–5393.
+- Liao and Vaughan (2024) Q Vera Liao and Jennifer Wortman Vaughan.
+2024.
+AI Transparency in the Age of LLMs: A Human-Centered Research Roadmap.
+Harvard Data Science Review (2024), 1–53.
+https://hdsr.mitpress.mit.edu/pub/aelql9qy/release/2
 - Madaio et al.
@@
 Administration and Policy in Mental Health and Mental Health Services Research 42 (2015), 533–544.
+https://pmc.ncbi.nlm.nih.gov/articles/PMC4012002/
 - Papakyriakopoulos et al.
@@
 An Empirical Study.
-(2024).
+Proceedings of the 32nd IEEE/ACM International Conference on Program Comprehension (2024).
+https://mdipenta.github.io/files/icpc2024.pdf
 - Phillips-Wren and Adya (2020) Gloria Phillips-Wren and Monica Adya.
@@
 System-Level Transparency of Machine Learning.
-(February 22 2022).
 https://ai.meta.com/research/publications/system-level-transparency-of-machine-learning/
@@
 http://arxiv.org/abs/2204.01075
+- Qi et al.
+(2023) Xiangyu Qi, Yi Zeng, Tinghao Xie, Pin-Yu Chen, Ruoxi Jia, Prateek Mittal, and Peter Henderson.
+2023.
+Fine-Tuning Aligned Language Models Compromises Safety, Even When Users Do Not Intend To!
+arXiv (October 5 2023).
+http://arxiv.org/abs/2310.03693
 - Raji and Yang (2020) Inioluwa Deborah Raji and Jingying Yang.
@@
 http://arxiv.org/abs/1912.06166
+- Rakova et al.
+(2021) Bogdana Rakova, Jingying Yang, Henriette Cramer, and Rumman Chowdhury.
+2021.
+Where responsible AI meets reality: Practitioner perspectives on enablers for shifting organizational practices.
+Proceedings of the ACM on Human-Computer Interaction 5, CSCW1 (2021), 1–23.
+https://dl.acm.org/doi/10.1145/3449081
 - Reid and Williams (2023) Kathy Reid and Elizabeth T.
@@
 http://arxiv.org/abs/2202.13028
+- Sambasivan et al.
+(2021) Nithya Sambasivan, Shivani Kapania, Hannah Highfill, Diana Akrong, Praveen Paritosh, and Lora M Aroyo.
+2021.
+“Everyone wants to do the model work, not the data work”: Data Cascades in High-Stakes AI.
+In proceedings of the 2021 CHI Conference on Human Factors in Computing Systems.
+1–15.
+https://dl.acm.org/doi/abs/10.1145/3411764.3445518
 - Shen et al.
@@
 https://doi.org/10.1145/3351095.3372870
-- Srinivasan et al. (2021) Ramya Srinivasan, Emily Denton, Jordan Famularo, Negar Rostamzadeh, Fernando Diaz, and Beth Coleman. 2021. Artsheets for Art Datasets.
+- Solaiman et al.
+(2023) Irene Solaiman, Zeerak Talat, William Agnew, Lama Ahmad, Dylan Baker, Su Lin Blodgett, Canyu Chen, Hal Daumé III, Jesse Dodge, Isabella Duan, et al.
+2023.
+Evaluating the Social Impact of Generative AI Systems in Systems and Society.
+arXiv preprint arXiv:2306.05949 (2023).
+https://arxiv.org/abs/2306.05949
+- Srinivasan et al.
+(2021) Ramya Srinivasan, Emily Denton, Jordan Famularo, Negar Rostamzadeh, Fernando Diaz, and Beth Coleman.
+2021.
+Artsheets for Art Datasets.
+In Proceedings of the Neural Information Processing Systems Track on Datasets and Benchmarks 1 (NeurIPS Datasets and Benchmarks 2021).
+https://datasets-benchmarks-proceedings.neurips.cc/paper/2021/hash/9b8619251a19057cff70779273e95aa6-Abstract-round2.html
 - Stoyanovich and Howe (2019) Julia Stoyanovich and Bill Howe. 2019. Nutritional Labels for Data and Models. Bulletin of the IEEE Computer Society Technical Committee on Data Engineering (2019).
@@
 https://doi.org/10.1145/3357384.3357853
+- Suresh et al.
+(2021) Harini Suresh, Steven R Gomez, Kevin K Nam, and Arvind Satyanarayan.
+2021.
+Beyond expertise and roles: A Framework to Characterize the Stakeholders of Interpretable Machine Learning and Their Needs.
+In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems.
+1–16.
+https://dl.acm.org/doi/10.1145/3411764.3445088
+- Suri (2011) Harsh Suri. 2011. Purposeful sampling in qualitative research synthesis. Qualitative research journal 11, 2 (2011), 63–75.
 - Tagliabue et al. (2021) Jacopo Tagliabue, Ville Tuulos, Ciro Greco, and Valay Dave. 2021. DAG Card Is the New Model Card. arXiv (November 20 2021). http://arxiv.org/abs/2110.13601
@@
 https://doi.org/10.48550/arXiv.1706.03762
+- Weidinger et al.
+(2023) Laura Weidinger, Maribeth Rauh, Nahema Marchal, Arianna Manzini, Lisa Anne Hendricks, Juan Mateos-Garcia, Stevie Bergman, Jackie Kay, Conor Griffin, Ben Bariach, et al.
+2023.
+Sociotechnical Safety Evaluation of Generative AI Systems.
+arXiv preprint arXiv:2310.11986 (2023).
+https://arxiv.org/abs/2310.11986
 - Winecoff and Watkins (2022) Amy A.
@@
 https://doi.org/10.1145/3183713.3193568
+- Yurrita et al.
+(2023) Mireia Yurrita, Agathe Balayn, and Ujwal Gadiraju.
+2023.
+Generating Process-Centric Explanations to Enable Contestability in Algorithmic Decision-Making: Challenges and Opportunities.
+arXiv preprint arXiv:2305.00739 (2023).
+https://arxiv.org/abs/2305.00739
 - Zheng et al.